### v1

# Learning under Distributional Drift: Reproducibility as an Intrinsic Statistical Resource

Abstract
Statistical learning under distributional drift remains insufficiently characterized: when each observation alters the data-generating law, classical generalization bounds can collapse. We introduce a new statistical primitive, the reproducibility budget $C_{T}$, which quantifies a system’s finite capacity for statistical reproducibility—the extent to which its sampling process can remain governed by a consistent underlying distribution in the presence of both exogenous change and endogenous feedback. Formally, $C_{T}$ is defined as the cumulative Fisher–Rao path length of the coupled learner–environment evolution, measuring the total distributional motion accumulated during learning. From this construct we derive a drift–feedback generalization bound of order $O(T^{-1/2}+C_{T}/T)$, and we prove a matching minimax lower bound showing that this rate is minimax-optimal. Consequently, the results establish a reproducibility speed limit: no algorithm can achieve smaller worst–case generalization error than that imposed by the average Fisher–Rao drift rate $C_{T}/T$ of the data-generating process. The framework situates exogenous drift, adaptive data analysis, and performative prediction within a common geometric structure, with $C_{T}$ emerging as the intrinsic quantity measuring distributional motion across these settings.
Keywords: information geometry, statistical learning theory, non-stationary learning, generalization bounds, distributional shift
1 Introduction
Modern learning systems often operate in self-modifying environments, where learning alters the very distribution it learns from. A recommender policy reshapes user preferences as it updates (Perdomo2020); adaptive experiments change the distribution from which subsequent data are drawn (li2024); and reinforcement-driven agents modify the state transitions that determine future feedback (Sutton1998; Dar2003; AstromMurray2008). In all such cases, the learner is not a passive observer of a stationary process but an active participant in an evolving distributional flow. This feedback coupling breaks not only stationarity but the assumption of statistical reproducibility—that future data could, in principle, be redrawn from the same law (Vapnik1998).
1.1 Motivation
Given samples $\{(x_{i},y_{i})\}_{i=1}^{T}$ drawn i.i.d. from a fixed distribution $p(x,y)$, population and empirical risks coincide asymptotically, with convergence rate $O(T^{-1/2})$ (Vapnik1998). Once the data-generating process itself changes, these guarantees collapse. When distributional shift depends on the learner’s trajectory, the change is no longer exogenous but endogenous: a function of the learner’s own actions and their cumulative impact on the environment. In such settings, convergence is governed not by sample size, but by how far the underlying distribution moves over the course of learning, a quantity naturally measured in the statistical geometry of the model.
Consider an adaptive recommender with policy $\pi_{t}$ that selects content for users while updating its model from observed responses. Each recommendation depends on past data yet also shapes future data, since user preferences evolve with exposure. Let the environment be parameterized by $\theta_{t}$, governing $p_{\theta_{t}}(x,y)$, and evolve according to
$(x_{t},y_{t})\sim p_{\theta_{t}},\qquad\theta_{t+1}=F(\theta_{t},u_{t}),\qquad u_{t}\sim\pi_{t}(\mathcal{F}_{t}),$ (1)
where $\mathcal{F}_{t}$ is the filtration of all past observations and $u_{t}$ is some action chosen by the learner. Even if $F$ is smooth and deterministic, its dependence on $u_{t}$ ensures that $\theta_{t+1}$ reflects the learner’s internal decisions, producing a coupled stochastic process with a natural Fisher–Rao geometry. The family $\{p_{\theta}\}$ forms a statistical manifold whose intrinsic Riemannian metric—the Fisher information—measures how sensitively the distribution changes under infinitesimal perturbations of $\theta$. Thus the learner–environment interaction induces geometric motion in Fisher–Rao distance. The resulting sequence $\{(x_{t},y_{t})\}$ is non-reproducible: rerunning the same algorithm would yield a distinct trajectory unless the entire feedback loop were exactly replicated.
The same feedback-driven non-reproducibility arises in adaptive testing, model-in-the-loop simulation, and autonomous control. This raises a fundamental question: how far can the learner–environment system move before classical guarantees break? To answer this, we develop a geometric framework for quantifying distributional motion during learning. Across these closed-loop regimes, we analyze this coupling at the level of statistical generalization, showing that diverse non-stationary settings emerge as special cases of a single geometric bound governed by a finite reproducibility budget.
Figure 1: Coupled drift process. The learner’s policy $\pi_{t}$ and exogenous influence $\eta_{t}$ act on the environment, evolving $\theta_{t}$ to $\theta_{t+1}$ under $F(\theta_{t},u_{t},\eta_{t})$. This new state defines the next data distribution $p_{\theta_{t+1}}$, closing the feedback loop. Exogenous factors $\eta_{t}$ perturb $\theta_{t}$ externally, while endogenous feedback arises from the learner’s own actions.
1.2 Conceptual View
We view the data-generating process as a trajectory $\{\theta_{t}\}$ on a statistical manifold, whose discrete path length represents cumulative information change (AmariNagaoka2000). Any smooth divergence admits a second-order expansion whose Hessian defines a local Riemannian metric (Amari2016). For the broad class of $f$-divergences, including KL, $\chi^{2}$, Hellinger, and $\alpha$-divergences, this induced metric coincides with the Fisher–Rao geometry. Fisher–Rao is therefore the canonical invariant Riemannian metric on a statistical model (cencov1982; AmariNagaoka2000). It is preserved under smooth reparameterizations and captures displacement intrinsic to the model family rather than to any particular coordinate system. Although different divergences quantify global discrepancies between distributions, their local quadratic structure collapses to the same information metric, and only the Fisher metric enjoys this invariance property (cencov1982). At an infinitesimal scale, Fisher information provides the quadratic form that linearizes KL divergence and characterizes the distinguishability of nearby models. This is not a peculiarity of KL, but a reflection of the shared second-order geometry of all $f$-divergences. Consequently, Fisher–Rao arc length offers a natural intrinsic scale for describing how small distributional perturbations alter the statistical environment of a learner.
Accumulating these infinitesimal displacements yields an intrinsic Riemannian arc length (AmariNagaoka2000). This differs fundamentally from summing TV, Wasserstein, or KL increments, which are extrinsic divergences and do not compose additively along a path. What the arc length measures is not merely “how far the distribution moved” in a chosen divergence, but the total statistical distinguishability expended along the trajectory in the unique geometry that governs estimation. Generalization error then depends not only on the number of samples but also on the total path length traversed by the data distribution during learning.
Definition 1 (Statistical Reproducibility)
Reproducibility is a statistical property of a learning process: the extent to which repeated executions of the same algorithm would sample from comparable data-generating laws. Let $p_{\theta_{t}}(x,y)$ denote the distribution at time $t$ and $g_{\theta}$ the Fisher–Rao metric on $\Theta$. The reproducibility budget
$C_{T}\;=\;\sum_{t=1}^{T}\|\Delta\theta_{t}\|_{g_{\theta_{t}}}$
quantifies the cumulative Fisher–Rao path length of the trajectory $\{\theta_{t}\}$ and hence the total loss of reproducibility incurred during learning. The term “budget” emphasizes that reproducibility is finite. Each Fisher–Rao displacement consumes part of the available geometric capacity for drawing comparable samples. When $C_{T}$ grows large, reproducibility is exhausted.
Although the quantity $C_{T}$ is defined as the Fisher–Rao path length of the trajectory, its role in generalization is not assumed. Rather, the dependence of the risk on $C_{T}$ follows from the Fisher–Rao geometry and is shown to be minimax-optimal in Theorem 15.
While $C_{T}$ measures total Fisher–Rao distance, the motion of $\theta_{t}$ may arise from distinct causes. Some displacement is exogenous, driven by external change. The rest is endogenous, reflecting how the learner’s own policy alters the data-generating law. Decomposing each step’s Fisher length into these components gives
$C_{T}=\sum_{t=1}^{T}(d_{t}+\alpha\,\kappa_{t}^{(\mathcal{M})}),$ (2)
where $d_{t}$ captures environment-driven motion, $\kappa_{t}^{(\mathcal{M})}$ captures policy-sensitive motion, and $\alpha$ balances their contribution. The budget $C_{T}$ is intrinsic—the Fisher–Rao path length of the trajectory. The decomposition into $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$ is carried out entirely within the same Riemannian geometry, where both terms are derived from the Fisher metric and together approximate $\|\Delta\theta_{t}\|_{g_{\theta_{t}}}$. What is model-dependent is only which directions of the tangent space are activated by the environment versus the learner, not the geometric structure of the decomposition itself.
Our results show that the expected generalization deviation satisfies
$\mathbb{E}\,|\widehat{R}_{T}-R_{T}|\leq C_{0}T^{-1/2}+C_{1}T^{-1}\!\sum_{t}d_{t}+C_{2}T^{-1}\!\sum_{t}\kappa_{t}^{(\mathcal{M})},$ (3)
and that no algorithm operating within this drift–feedback model can improve upon the asymptotic rate $\Theta(T^{-1/2}+C_{T}/T)$. This defines a reproducibility speed limit: no learner can achieve smaller worst-case generalization error than that imposed by the average Fisher–Rao drift rate $C_{T}/T$ of the data–generating process. Earlier analyses bounded distributional change through variation or information budgets. Those quantities are extrinsic and problem-specific. The reproducibility budget $C_{T}$ is intrinsic. It is defined directly on the statistical manifold and admits a matching lower bound that prior formulations lack. It provides a geometric quantity that can be measured along the data-generating trajectory rather than imposed as an external regularity condition. A visual illustration is provided in Figure 2.
(a) Local drift decomposition in the tangent space $T_{\theta_{t}}\Theta$.
(b) Global Fisher–Rao path and cumulative length $C_{T}$.
Figure 2: Geometric intuition for learning under drift: (a) at each step the environment parameter $\theta_{t}$ lies on the statistical manifold $(\Theta,g_{\theta})$ and evolves via a tangent update $v_{t}\in T_{\theta_{t}}\Theta$, which decomposes into an exogenous component $d_{t}$ capturing distributional motion independent of the learner and an endogenous component $\kappa_{t}^{(\mathcal{M})}$ induced by learner feedback; the exponential map $\exp_{\theta_{t}}(v_{t})$ lifts this linearized update back onto the curved manifold, producing $\theta_{t+1}$ and realizing the first–order Fisher–geometric drift decomposition used in the analysis; (b) iterating these local movements generates a global Fisher–Rao trajectory $\theta_{0}\!\to\!\theta_{1}\!\to\!\cdots\!\to\!\theta_{T}$, whose intrinsic length $C_{T}\approx\sum_{t=0}^{T-1}d_{F}(\theta_{t+1},\theta_{t})$ measures the total statistical motion of the data–generating process. This cumulative Fisher path is the reproducibility budget that governs drift–induced generalization behavior.
1.3 Relation to Prior Frameworks
Several established frameworks have analyzed fragments of the broader picture studied here. While each captures an important aspect of instability, they are typically treated separately. Our work provides a geometric viewpoint that connects them.
Stationary learning (Vapnik1998; shalev2014) assumes fixed $p(x,y)$ and produces the familiar $O(T^{-1/2})$ convergence guarantees derived from concentration under i.i.d. sampling. Nonstationary stochastic optimization (Besbes2015; Zhao2024) introduces a variation budget that measures how much the target distribution can change exogenously over time. The metric of variation is typically a total-variation or Wasserstein norm, which constrains external motion but does not model coupling to the learner’s policy. Related approaches bound performance under drifting distributions using stability or Rademacher techniques (Blanchard2021; kuznetsov16; Mohri2012) but likewise do not explicitly model feedback between learner and environment. Performative prediction (Perdomo2020) does explicitly introduce feedback between model and environment, focusing on equilibrium conditions where that feedback vanishes ($p_{\theta_{t+1}}\!\approx\!p_{\theta_{t}}$). Follow-up work on performative prediction has refined aspects of this coupling (izzo21; jagadeesan22), but continues to analyze convergence to fixed points rather than the explicit accumulation of temporal drift. Adaptive data analysis (Dwork2015; rogers20; bassily2016) addresses feedback in the statistical query setting and bounds overfitting through mutual-information or stability arguments, yet does so within a static population that does not itself evolve (see also russo2016; Seldin2012; AlquierGuedj2018).
These approaches define four distinct ways to measure instability: sample variance, variation budgets, equilibrium deviation, and information flow. The present work connects them by treating all forms of drift as motion on a statistical manifold. The total motion of the data distribution decomposes as
$\underbrace{d_{t}}_{\text{exogenous drift}}\quad+\quad\underbrace{\kappa_{t}^{(\mathcal{M})}}_{\text{endogenous (policy-sensitive) drift}},$ (4)
where $d_{t}$ measures environment-driven change and $\kappa_{t}^{(\mathcal{M})}$ measures learner-induced motion under the Fisher–Rao metric $g_{\theta}$ on $\Theta$. Although the decomposition in 4 is simple to state, its validity within the Fisher–Rao geometry is not obvious a priori. This formulation shows that the classical regimes can be interpreted as limiting cases distinguished by which source of motion dominates.
Table 1: Learning regimes recovered as limits of the drift–feedback bound. Each corresponds to nullifying either the exogenous drift $d_{t}$ or the policy-sensitivity term $\kappa_{t}^{(\mathcal{M})}$.
Learning regime Active terms Representative work Notes
Stationary i.i.d. $\sqrt{T}$ term only Vapnik1998 No drift or feedback; fixed $p(x,y)$.
Exogenous drift $d_{t}$ term Besbes2015 Distribution changes externally over time.
Performative equilibrium $d_{t}\!\to\!0,\ \kappa_{t}^{(\mathcal{M})}\!\to\!0$ Perdomo2020 Coupled learner–environment system converging to a performative equilibrium.
Adaptive data analysis $\kappa_{t}^{(\mathcal{M})}$ term Dwork2015 Purely feedback-driven nonreproducibility.
A broader literature also bounds performance in nonstationary and online optimization through dynamic regret or excess-risk budgets (Zhao2024), while information-theoretic analyses of adaptive learning express related limits in terms of mutual-information accumulation between data and model parameters (russo2016; xu2017). In statistical physics, learning has been interpreted as a thermodynamic process in which information gain is constrained by entropy production (Still2012; Goldt2017; ortega2013thermodynamics). All of these perspectives describe some notion of limited information flow, but they do so with external metrics or equilibrium arguments rather than an intrinsic geometric invariant.
The reproducibility budget $C_{T}$ provides such an invariant. It measures the total Fisher–Rao path length of the coupled learner–environment trajectory, bringing variation budgets, information-based stability, and thermodynamic perspectives into a common geometric viewpoint. Classical $O(T^{-1/2})$ generalization corresponds to the zero-motion limit, while the $O(C_{T}/T)$ term quantifies the finite amount of information geometry that can be traversed before reproducibility fails. In this sense, $C_{T}$ places the separate notions of variation, feedback, and information accumulation within a single statistical geometry of drift.
1.4 Contributions and Significance
The reproducibility budget $C_{T}$ serves as a new statistical primitive that quantifies the intrinsic information–geometric distance traversed by a learning process. Where existing theories bound isolated sources of instability (exogenous variation or adaptive feedback), $C_{T}$ provides a geometric invariant that governs both. It reframes reproducibility not as an assumption but as a finite resource whose consumption limits generalization under self–altering dynamics. Building on this primitive, we develop a unified theory of learning under distributional drift in which environmental change and learner feedback jointly shape the data–generating process. The resulting analysis extends stationary learning theory to closed–loop systems and establishes a quantitative limit on how rapidly a learner can adapt without violating reproducibility.
The main contributions are:
  1. 1.
    We develop a unified closed-loop framework for learning under endogenous drift, modeling the data-generating process as a trajectory on a statistical manifold.
  2. 2.
    We derive a geometric decomposition of distributional motion into exogenous and policy-sensitive components $(d_{t},\kappa_{t}^{(\mathcal{M})})$, yielding a tractable reproducibility budget $C_{T}$.
  3. 3.
    We establish finite-sample drift–feedback generalization bounds that separate sampling variability, exogenous drift, and policy-induced feedback.
  4. 4.
    We prove a matching minimax lower bound, establishing a reproducibility speed limit and demonstrating the necessity of the policy-sensitivity term.
  5. 5.
    We provide an information–geometric interpretation connecting cumulative drift to Fisher–Rao arc length and local KL structure, unifying variation, stability, and information budgets under a single intrinsic geometry.
  6. 6.
    We validate the theory in controlled linear–Gaussian systems, a nonlinear teacher–learner experiment, and a metric-alignment study, confirming the additive and predictive structure of the reproducibility budget.
The resulting framework is conceptually simple, but identifying a geometric decomposition of drift into exogenous and policy-sensitive components is not automatic. It is not obvious a priori that such a decomposition should exist or behave additively within the intrinsic Fisher–Rao geometry.
The remainder of the paper is organized as follows. Section 2 introduces preliminary tools that will be used in the subsequent analysis; Section 3 formalizes the endogenously drifting process and its metric structure; Section 4 defines the drift measures and reproducibility budget; Section 5 presents the main theorems and lower bounds; Section 6 develops the information–geometric interpretation; Section 7 reports empirical validation; and Section 8 discusses implications and extensions.
2 Preliminaries and Lemmas
This section collects the technical tools used throughout the paper. We formalize the probabilistic and geometric setting, state regularity assumptions ensuring a well-defined Fisher geometry, and record several auxiliary inequalities connecting concentration, transport, and path length on the statistical manifold $(\Theta,g_{\theta})$. All results hold under the standing assumptions below and are invoked repeatedly in later sections. Full proofs of all nonstandard or technically involved statements in this section are deferred to Appendix A.
2.1 Measure–Metric Foundations
We begin by formalizing the geometric structure induced by the parametric family $\{p_{\theta}\}_{\theta\in\Theta}$. This structure provides the notion of “distance” between nearby distributions and will be used throughout to quantify the cumulative drift of the data-generating process.
Let $\mathcal{H}$ denote the learner’s hypothesis class, and $f\in\mathcal{H}$ a model mapping $\mathcal{X}\to\mathcal{Y}$ with measurable loss $\ell(f(x),y)$ that is $\sigma$–sub-Gaussian under each $p_{\theta}$. Let $(\mathcal{X},\mathcal{Y})$ denote the input–output space with the product $\sigma$–algebra, and let $\Theta\subset\mathbb{R}^{d}$ be a smooth parameter manifold indexing a family of data-generating distributions $\{p_{\theta}(x,y)\}_{\theta\in\Theta}$. Expectations with respect to $p_{\theta}$ are written $\mathbb{E}_{\theta}[\cdot]$.
Definition 2 (Fisher–Rao Metric and Distance)
For each $\theta\in\Theta$, let $T_{\theta}\Theta$ denote the tangent space of $\Theta$ at $\theta$. We write $\langle\cdot,\cdot\rangle$ for the standard Euclidean inner product on $\mathbb{R}^{d}$ and $\|\cdot\|$ for the associated norm.
The Fisher–Rao metric is the Riemannian metric $g_{\theta}:T_{\theta}\Theta\times T_{\theta}\Theta\to\mathbb{R}$ defined by
$g_{\theta}(v,w)\penalty 10000\ =\penalty 10000\ \mathbb{E}_{\theta}\!\left[\langle\nabla_{\theta}\log p_{\theta}(x,y),v\rangle\,\langle\nabla_{\theta}\log p_{\theta}(x,y),w\rangle\right],\qquad v,w\in T_{\theta}\Theta,$
where $\nabla_{\theta}$ denotes the gradient with respect to the parameter $\theta$. In local coordinates $\theta=(\theta^{1},\dots,\theta^{d})$, we write $\partial_{i}=\partial/\partial\theta^{i}$ and obtain the metric tensor entries
$g_{ij}(\theta)\penalty 10000\ =\penalty 10000\ \mathbb{E}_{\theta}\!\left[\partial_{i}\log p_{\theta}(x,y)\,\partial_{j}\log p_{\theta}(x,y)\right].$
For a tangent vector $v\in T_{\theta}\Theta$, the norm induced by $g_{\theta}$ is
$\|v\|_{g_{\theta}}\penalty 10000\ =\penalty 10000\ \sqrt{g_{\theta}(v,v)}\penalty 10000\ =\penalty 10000\ \sqrt{v^{\top}g(\theta)\,v},$
so that a differentiable curve $\theta:[0,1]\to\Theta$ has infinitesimal squared line element
$ds^{2}\penalty 10000\ =\penalty 10000\ \dot{\theta}^{\top}g(\theta)\,\dot{\theta},$
where $\dot{\theta}$ is the time derivative of $\theta$.
The associated Fisher–Rao geodesic distance between $\theta,\theta^{\prime}\in\Theta$ is
$d_{F}(\theta,\theta^{\prime})\penalty 10000\ =\penalty 10000\ \inf_{\gamma\in\Gamma(\theta,\theta^{\prime})}\int_{0}^{1}\|\dot{\gamma}(s)\|_{g_{\gamma(s)}}\,ds,$
[…]
Several equivalent formulations of the Fisher–Rao geometry will be used interchangeably. They connect score covariance, local KL structure, and geodesic motion on $\Theta$.
Lemma 3 (Equivalent Characterizations of the Fisher–Rao Metric)
Assume the parametric family $\{p_{\theta}\}_{\theta\in\Theta}$ is regular in the sense that $\theta\mapsto\log p_{\theta}(x,y)$ is twice continuously differentiable, differentiation and integration may be interchanged, and the Fisher information matrix $g(\theta)$ is finite and positive definite on an open neighborhood of $K$. Then the Fisher–Rao metric of Definition 2 admits the following equivalent characterization:
  1. (i)
    Score covariance (coordinate form). In any local coordinate chart,
    $g_{ij}(\theta)=\mathbb{E}_{\theta}\!\left[\partial_{i}\log p_{\theta}(x,y)\,\partial_{j}\log p_{\theta}(x,y)\right],$
    so the Fisher information matrix is the covariance matrix of the score vector $s_{\theta}=\nabla_{\theta}\log p_{\theta}(x,y)$.
  2. (ii)
    Local KL expansion. For the Kullback–Leibler divergence $D_{\mathrm{KL}}(p\|q)=\mathbb{E}_{p}[\log(p/q)]$,
    $D_{\mathrm{KL}}(p_{\theta+\Delta\theta}\,\|\,p_{\theta})\penalty 10000\ =\penalty 10000\ \tfrac{1}{2}\,\|\Delta\theta\|_{g_{\theta}}^{2}\penalty 10000\ +\penalty 10000\ o\!\left(\|\Delta\theta\|_{g_{\theta}}^{2}\right),\qquad\Delta\theta\to 0.$
    Thus the Fisher norm is the infinitesimal square root of local KL divergence.
These identities require only the standard regularity conditions on $\{p_{\theta}\}$ stated explicitly below. They provide the geometric tools used throughout the analysis.
2.2 Regularity and Curvature Bounds
The analysis throughout relies on standard regularity properties of the statistical model and loss. We impose the following standing assumptions. For clarity, we write $R(\theta,f)=\mathbb{E}_{(x,y)\sim p_{\theta}}[\ell(f(x),y)]$ for the population risk of a fixed model $f$ under distribution $p_{\theta}$. When $\theta=\theta_{t}$ evolves over time, this coincides with the time-indexed population quantities $R_{t}$ introduced later.
Assumption 1 (Local Statistical Regularity)
The realized trajectory $\{\theta_{t}\}_{t=1}^{T}$ remains inside a compact region $K\subset\Theta$ on which the following local regularity conditions hold.
  1. 1.
    Finite Fisher information. For every $\theta\in K$, the Fisher information matrix $g(\theta)$ is finite, positive definite, and varies smoothly on $K$. Equivalently,
    $\sup_{t}\mathbb{E}_{p_{\theta_{t}}}\bigl\|s_{\theta_{t}}(x,y)\bigr\|^{2}<\infty,$
    where $s_{\theta}(x,y)=\nabla_{\theta}\log p_{\theta}(x,y)$ is the score.
  2. 2.
    Local smoothness of the model. The log-density $\theta\mapsto\log p_{\theta}(x,y)$ is twice continuously differentiable in a neighborhood of $K$, and differentiation and integration interchange uniformly on $K$.
  3. 3.
    Local bounded geometry. Since $g$ is smooth on the compact set $K$, its Christoffel symbols and curvature tensor are uniformly bounded on $K$. In normal coordinates around any $\theta\in K$,
    $g_{ij}(x)=\delta_{ij}+O(\|x\|^{2}),$
    and the $O(\|x\|^{2})$ terms are controlled uniformly over $\theta\in K$. These bounds are used only to control the quadratic remainder in Lemma 4.
  4. 4.
    Sub-Gaussian sampling noise. The martingale-difference noise in the empirical risk satisfies the conditional sub-Gaussian mgf bound with variance proxy $\sigma^{2}$:
    $\mathbb{E}\!\left[\exp(\lambda Z_{t})\mid\mathcal{F}_{t-1}\right]\leq\exp(\tfrac{1}{2}\sigma^{2}\lambda^{2}).$
Assumption 2 (Local Dynamics Regularity)
  1. 1.
    Finite-energy controls. The learner’s controls satisfy
    $\mathbb{E}\|u_{t}\|^{2}\;\leq\;B,\qquad t=1,\dots,T.$
    This ensures that control-induced perturbations remain within a region where the linearization below is valid.
  2. 2.
    Local differentiability in the control variable. For each $\theta\in K$ and each realized $\eta_{t}$, the map $u\mapsto F(\theta,u,\eta_{t})$ is differentiable at $u=0$, and admits the local expansion
    $F(\theta_{t},u_{t},\eta_{t})=F(\theta_{t},0,\eta_{t})+J_{u}F(\theta_{t},0,\eta_{t})\,u_{t}+r_{t},$
    with a quadratic remainder
    $\|r_{t}\|_{g_{\theta_{t}}}\;\leq\;c\,\|u_{t}\|^{2},$
    for a constant $c$ uniform over $\theta_{t}\in K$ and all admissible $u_{t}$.
  3. 3.
    Trajectory confinement. The update rule $\theta_{t+1}=F(\theta_{t},u_{t},\eta_{t})$ keeps the trajectory inside the regularity region $K$, so that the statistical and geometric bounds in Assumption 1 remain valid for all $t$.
These assumptions hold only along the realized trajectory and are used to control the local geometric terms that enter the proofs. All constants are understood to apply on the region visited by $\{\theta_{t}\}$. They provide a standard setting in which the Fisher geometry and concentration bounds are well behaved. In Section 7 we show that the same additive structure remains predictive in nonlinear neural environments where the model family is misspecified and the regularity conditions can hold, at best, locally along the path traced by training.
Under the curvature bounds in Assumption 1, geodesic distance is locally equivalent to the Fisher norm. Lemma 4 makes this precise.
Lemma 4 (Local geodesic equivalence)
Let $(\Theta,g)$ be a statistical manifold equipped with the Fisher metric. Fix $\theta\in\Theta$ and suppose the sectional curvature is bounded in a neighborhood of $\theta$. Then there exists $r>0$ and a constant $C>0$ such that for all $\Delta\theta$ with $\|\Delta\theta\|_{g_{\theta}}\leq r$,
$d_{F}(\theta,\theta+\Delta\theta)=\|\Delta\theta\|_{g_{\theta}}\,\bigl(1+E(\Delta\theta)\bigr),\qquad|E(\Delta\theta)|\leq C\,\|\Delta\theta\|_{g_{\theta}}.$ (5)
In particular,
$(1-Cr)\,\|\Delta\theta\|_{g_{\theta}}\;\leq\;d_{F}(\theta,\theta+\Delta\theta)\;\leq\;(1+Cr)\,\|\Delta\theta\|_{g_{\theta}}.$
Proof [Proof sketch] In normal coordinates at $\theta$, the Fisher metric satisfies $g_{ij}(x)=\delta_{ij}+O(\|x-\theta\|^{2})$, and the exponential map obeys $\operatorname{Exp}_{\theta}(v)=\theta+v+O(\|v\|^{2})$ (AmariNagaoka2000; Lee97). Thus if $\theta^{\prime}=\theta+\Delta\theta$ is sufficiently close to $\theta$, inverting the latter relation yields $v=\Delta\theta+O(\|\Delta\theta\|^{2})$. Gauss’s lemma then gives $d_{F}(\theta,\theta^{\prime})=\|v\|_{g_{\theta}}=\|\Delta\theta\|_{g_{\theta}}+O(\|\Delta\theta\|_{g_{\theta}}^{2})$, and shrinking the neighborhood of $\theta$ yields the stated uniform two–sided inequality.
2.3 Concentration Tools
We use standard concentration inequalities to control sampling variability and relate distributional motion to changes in expected loss.
Let $\{Z_{t}\}_{t=1}^{T}$ be a martingale difference sequence adapted to $\{\mathcal{F}_{t}\}$ and assume each $Z_{t}$ is conditionally $\sigma_{t}$–sub-Gaussian, i.e.
$\mathbb{E}\!\left[\exp\!\big(\lambda Z_{t}\big)\mid\mathcal{F}_{t-1}\right]\penalty 10000\ \leq\penalty 10000\ \exp\!\left(\tfrac{1}{2}\sigma_{t}^{2}\lambda^{2}\right)\qquad\text{for all }\lambda\in\mathbb{R}.$
Set $S_{T}=\sum_{t=1}^{T}Z_{t}$ and $V_{T}=\sum_{t=1}^{T}\sigma_{t}^{2}$.
Lemma 5 (Sub-Gaussian martingale deviation)
For any $\eta>0$,
$\Pr\!\big(|S_{T}|\geq\eta\big)\penalty 10000\ \leq\penalty 10000\ 2\exp\!\left(-\frac{\eta^{2}}{2V_{T}}\right).$
This is a standard sub-Gaussian martingale concentration argument (Boucheron2013; bercu2015).
Corollary 6 (Expected deviation)
Under the conditions of Lemma 5,
$\mathbb{E}\,|S_{T}|\penalty 10000\ \leq\penalty 10000\ \sqrt{2\pi\,V_{T}},\qquad\mathbb{E}\!\left|\frac{1}{T}\sum_{t=1}^{T}Z_{t}\right|\penalty 10000\ \leq\penalty 10000\ \frac{\sqrt{2\pi\,V_{T}}}{T}.$
In particular, if each $Z_{t}$ is $\sigma$–sub-Gaussian, so that $V_{T}\leq\sigma^{2}T$, then
$\mathbb{E}\!\left|\frac{1}{T}\sum_{t=1}^{T}Z_{t}\right|\penalty 10000\ \leq\penalty 10000\ \frac{\sqrt{2\pi}\,\sigma}{\sqrt{T}}.$
Proof [Proof sketch] By Lemma 5, $\Pr(|S_{T}|\geq\eta)\leq 2\exp(-\eta^{2}/(2V_{T}))$. Integrating this tail bound,
$\mathbb{E}\,|S_{T}|\penalty 10000\ =\penalty 10000\ \int_{0}^{\infty}\Pr(|S_{T}|\geq\eta)\,d\eta\penalty 10000\ \leq\penalty 10000\ \int_{0}^{\infty}2\exp\!\Big(-\frac{\eta^{2}}{2V_{T}}\Big)\,d\eta\penalty 10000\ =\penalty 10000\ \sqrt{2\pi\,V_{T}}.$
Dividing by $T$ yields the averaged bound. Full details appear in Appendix A.
2.4 Length Convergence and Path Integrals
The reproducibility budget detailed later is a discrete Fisher path length. The following lemma identifies this discrete quantity with the continuous arc length of its canonical piecewise–geodesic interpolant. In the small–step regime, this provides the continuous representation to which the drift increments of Section 4.2 converge.
Lemma 7 (Discrete–Continuous Fisher–Rao Length Identity)
Let $\theta_{0},\ldots,\theta_{T}\in\Theta$ be a trajectory, and let $\gamma_{T}:[0,1]\to\Theta$ denote the piecewise minimizing Fisher–Rao geodesic satisfying $\gamma_{T}(t/T)=\theta_{t}$ for all $t$. Then
$\sum_{t=0}^{T-1}d_{F}(\theta_{t+1},\theta_{t})\;=\;\int_{0}^{1}\big\|\dot{\gamma}_{T}(s)\big\|_{g_{\gamma_{T}(s)}}\,ds.$ (6)
Proof Each segment $\gamma_{T}|_{[t/T,(t+1)/T]}$ is a constant–speed minimizing geodesic from $\theta_{t}$ to $\theta_{t+1}$, so its Riemannian length equals $d_{F}(\theta_{t+1},\theta_{t})$. Summing over segments gives the identity.
2.5 Summary of Constants
For reference, the principal constants are:
$\sigma$    sub-Gaussian loss parameter
$L_{p}$ Lipschitz constant linking risk to $W_{1}$
$G_{\max}$ bound on Fisher curvature (Christoffel symbols)
$c_{d}$ local geodesic equivalence constant
Together, Lemmas 47 form the technical backbone used in Sections 35. They ensure that Fisher distances, path lengths, and concentration terms behave smoothly under Assumption 1, enabling the drift–feedback analysis developed next.
3 Problem Setup
Having established the geometric and concentration tools in Section 2, we now instantiate them in a stochastic process that captures learning under endogenous drift—a regime in which the learner’s own actions alter the distribution from which future observations are drawn. Our goal is to express this closed-loop evolution directly in the Fisher geometry defined earlier, so that distributional motion can be quantified in intrinsic terms.
We begin by specifying the coupled learner–environment dynamics. Let $(\mathcal{X},\mathcal{Y})$ denote the input–output space and $\Theta$ the parameter manifold indexing the family of distributions $\{p_{\theta}(x,y)\}_{\theta\in\Theta}$. At each discrete time $t\in\{1,\ldots,T\}$,
$(x_{t},y_{t})\sim p_{\theta_{t}},\qquad\theta_{t+1}=F(\theta_{t},u_{t},\eta_{t}),\qquad u_{t}\sim\pi_{t}(\mathcal{F}_{t}),$ (7)
where $\mathcal{F}_{t}=\sigma(x_{1},y_{1},\ldots,x_{t},y_{t})$ is the natural filtration. The map $F$ governs the environment’s evolution, $\pi_{t}$ is a (possibly history–dependent) policy, and $\eta_{t}$ represents exogenous latent influences.
Definition 8 (Endogenously Drifting Process)
A learning process $\{(x_{t},y_{t},\theta_{t},u_{t})\}_{t=1}^{T}$ satisfies endogenous drift if the transition $F(\theta_{t},u_{t},\eta_{t})$ depends on the control $u_{t}$, thereby coupling the evolution of $\theta_{t}$ to the policy $\pi_{t}$. The variable $\eta_{t}$ represents the exogenous influence acting on the environment: a collection of factors, deterministic or stochastic, that evolve independently of $u_{t}$ and contribute exogenous drift.
Example 1 (Linear–Gaussian environment)
A classical linear instance of (7) is
$F(\theta_{t},u_{t},\eta_{t})\penalty 10000\ =\penalty 10000\ \theta_{t}+Au_{t}+B\eta_{t},$
where the matrices $A$ and $B$ determine sensitivity to the learner’s action $u_{t}$ and to exogenous influences $\eta_{t}$ (AstromMurray2008). The term $Au_{t}$ captures policy-induced motion and vanishes in the purely exogenous case $A=0$, while the term $B\eta_{t}$ captures background evolution and vanishes when $B=0$, yielding drift driven entirely by the learner.
This instance satisfies the semantics of Definition 8 exactly and provides a setting in which Fisher path length, drift magnitudes $(d_{t},\kappa_{t}^{(\mathcal{M})})$, and the reproducibility budget $C_{T}$ can be computed in closed form. It also underlies the linear–Gaussian experiments in Section 7, where the impact of endogenous and exogenous drift can be isolated directly.
A canonical learning problem that instantiates this model is linear regression with Gaussian covariates and Gaussian noise. Let $x_{t}\sim\mathcal{N}(\mu_{t},\Sigma)$ and $y_{t}=w^{\top}x_{t}+\epsilon_{t}$ with $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$. Here the environment parameter is the drifting mean $\theta_{t}=\mu_{t}$, whose evolution is governed by the same linear–Gaussian update $F$. Drift in $\mu_{t}$ induces distributional drift in $p_{\theta_{t}}$, and the optimal predictor moves linearly with $\mu_{t}$, making this an exact instance of the linear–Gaussian framework above.
Endogenous drift thus produces a closed-loop system in which learner and environment co-evolve. If $F$ is independent of $u_{t}$, the setting reduces to non-stationary optimization with purely exogenous drift. If $F$ is constant, it reduces to the classical i.i.d. regime. The general case—where both influences are present—yields an evolving distributional flow driven jointly by background change and policy-induced perturbations.
A natural question is whether such coupling might guide the process toward a performative equilibrium, so that policy-induced drift $\kappa_{t}^{(\mathcal{M})}$ eventually vanishes. In general, this need not occur. The joint evolution of $(f_{t},\theta_{t})$ is not assumed to be a gradient flow on a shared potential, and the environment map $F$ need not be contractive. Even when fixed points exist, stochasticity, misspecification, or misaligned objectives can yield persistent cycles or wandering trajectories. Prior work on performative prediction typically imposes strong stability conditions to study equilibria (Perdomo2020). Our analysis instead addresses the generic case in which the learner–environment system continues to move along the statistical manifold and the cumulative motion grows with $T$.
With the dynamics specified, we now express the relevant statistical quantities along the evolving trajectory. The parameter space $\Theta$ carries the Fisher–Rao metric $g_{\theta}$ and geodesic distance $d_{F}$ introduced in Section 2. All norms, distances, and path lengths are understood with respect to this metric, with local equivalence and curvature bounds provided by the regularity conditions.
Our focus is the trajectory–averaged generalization gap $\widehat{R}_{T}-R_{T}$, the natural notion of generalization under closed-loop drift where both $(f_{t})$ and $(\theta_{t})$ evolve over time. Define
$\widehat{R}_{T}(f,\pi)=\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}(x_{t}),y_{t}),\qquad R_{T}(f,\pi)=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{(x,y)\sim p_{\theta_{t}}}\!\left[\ell(f_{t}(x),y)\right],$
and let
$\Delta_{T}(f,\pi)\penalty 10000\ =\penalty 10000\ \bigl|\widehat{R}_{T}(f,\pi)-R_{T}(f,\pi)\bigr|$
denote the trajectory–averaged generalization error. We analyze the expectation $\mathbb{E}\,\Delta_{T}(f,\pi)$, which captures generalization under distributional drift.
To describe the environment’s motion on $(\Theta,g_{\theta})$, we separate exogenous and policy-sensitive components. The exogenous drift magnitude is
$d_{t}=d_{F}(\bar{\theta}_{t+1},\bar{\theta}_{t}),\qquad\bar{\theta}_{t+1}=F(\bar{\theta}_{t},0,\eta_{t}),$
capturing environment evolution independent of the learner. Expectations over $\eta_{t}$ are absorbed into $d_{t}$, and we write $F(\theta_{t},u_{t})$ for the nominal update.
Policy-sensitive drift is measured by
$\kappa_{t}^{(\mathcal{M})}=\|J_{u}F(\theta_{t},0,\eta_{t})\,u_{t}\|_{g_{\theta_{t}}},$ (8)
the leading-order Fisher–metric motion induced by the learner’s action. Higher-order contributions will be handled in Section 4.
These components combine to form the reproducibility budget
$C_{T}\penalty 10000\ =\penalty 10000\ \sum_{t=1}^{T}\bigl(d_{t}+\alpha\,\kappa_{t}^{(\mathcal{M})}\bigr),$
a tractable surrogate for the Fisher path length of the trajectory. The scalar $\alpha>0$ is not a model parameter. It plays no role in the dynamics of $\theta_{t}$ and serves only to place $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$ on a common scale so they can be combined into the single quantity $C_{T}$.
Remark 9
The update (7) defines a trajectory on $(\Theta,g_{\theta})$. The quantities $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$ describe its incremental motion, while $C_{T}$ records the accumulated Fisher length.
Together, $d_{t}$, $\kappa_{t}^{(\mathcal{M})}$, and $C_{T}$ serve as the geometric primitives of the framework and anchor the drift–feedback bounds developed in Section 5.
4 Quantifying Drift and Feedback
The quantities introduced in Section 3 measure instantaneous motion on $(\Theta,g_{\theta})$ and describe how control and exogenous influences jointly move the data–generating process. We now develop how these local motions aggregate over time and how they connect to reproducibility and generalization.
4.1 Drift Decomposition
Because $F$ is smooth in the control variable (Assumption 2), the update admits the first–order expansion
$\Delta\theta_{t}\penalty 10000\ =\penalty 10000\ \underbrace{F(\theta_{t},0,\eta_{t})-\theta_{t}}_{\bar{\Delta}\theta_{t}}\;+\;\underbrace{J_{u}F(\theta_{t},0,\eta_{t})\,u_{t}}_{\text{policy term}}\;+\;r_{t},$ (9)
where $r_{t}$ collects curvature and higher–order terms. This separates environment–driven motion from the leading-order policy-sensitive contribution.
Conditioning on $\theta_{t}$ fixes the local Fisher metric, so the Taylor expansion and triangle inequality apply inside the conditional expectation. The exogenous and policy terms contribute linearly, while the curvature remainder contributes through the second–order term $\epsilon_{t}:=\mathbb{E}[\,\|r_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}\,]$. All expansions are taken in the ambient coordinate chart on $\Theta\subset\mathbb{R}^{d}$. The Fisher metric is applied only after the increment $\Delta\theta_{t}$ is computed. This gives
$\mathbb{E}\!\left[\|\Delta\theta_{t}\|_{g_{\theta_{t}}}\,\middle|\,\theta_{t}\right]\;\leq\;d_{t}+\alpha\,\kappa_{t}^{(\mathcal{M})}+\epsilon_{t},$ (10)
where $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$ are the drift magnitudes from Section 3, $\alpha$ balances their relative scale, and $\epsilon_{t}$ consists of second–order residuals with $\mathbb{E}[\epsilon_{t}]\penalty 10000\ =\penalty 10000\ O\!\left(\mathbb{E}[\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2}]\right)$, under the smoothness conditions on $F$ (AmariNagaoka2000; Absil2008). Equation (10) gives the basic link between first–order motion and its geometric decomposition. While the decomposition is algebraically simple, its validity in the intrinsic Fisher–Rao geometry is not automatic. It relies on the smoothness and curvature controls established in Section 2.
4.2 Cumulative Motion and the Reproducibility Budget
Accumulating these local steps along the trajectory yields the discrete Fisher path length
$\mathcal{A}_{\mathrm{path}}(T)\penalty 10000\ =\penalty 10000\ \sum_{t=1}^{T}\|\Delta\theta_{t}\|_{g_{\theta_{t}}}.$ (11)
Summing (10) over time,
$\mathbb{E}\,\mathcal{A}_{\mathrm{path}}(T)\penalty 10000\ \leq\penalty 10000\ C_{T}+\sum_{t=1}^{T}\epsilon_{t},$ (12)
where $C_{T}=\sum_{t=1}^{T}(d_{t}+\alpha\,\kappa_{t}^{(\mathcal{M})})$ is the reproducibility budget. When curvature residuals are small, $C_{T}$ approximates the Fisher arc length and acts as a tractable surrogate for the total statistical distance traveled.
Thus $d_{t}$ captures exogenous evolution, $\kappa_{t}^{(\mathcal{M})}$ captures policy-induced motion, and $C_{T}$ aggregates both into a single measure of cumulative information displacement.
4.3 From Geometric Motion to Risk Variation
We now relate geometric motion to risk. Lemma 11 gives
$|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})|\penalty 10000\ \leq\penalty 10000\ L_{p}\,d_{F}(\theta_{t+1},\theta_{t})\penalty 10000\ =\penalty 10000\ L_{p}\,\|\Delta\theta_{t}\|_{g_{\theta_{t}}}.$ (13)
Combining this with (10) yields
$\mathbb{E}\,|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})|\penalty 10000\ \leq\penalty 10000\ L_{p}\bigl(d_{t}+\alpha\,\kappa_{t}^{(\mathcal{M})}\bigr)+\tilde{\epsilon}_{t},$ (14)
with $\tilde{\epsilon}_{t}=L_{p}\,\epsilon_{t}$.
To connect these per-step population risk changes to the trajectory-averaged generalization gap, we first express the gap in a form that separates sampling deviations from the cumulative change in population risk.
Lemma 10 (Add–subtract expansion of the trajectory-averaged risk gap)
For any trajectory $\{(\theta_{t},f_{t},x_{t},y_{t})\}_{t=1}^{T}$, let
$Z_{t}:=\ell(f_{t}(x_{t}),y_{t})-R(\theta_{t},f_{t}),\qquad R_{t}:=R(\theta_{t},f_{t}).$
Then
$\bigl|\hat{R}_{T}-R_{T}\bigr|\;\leq\;\frac{1}{T}\Bigl|\sum_{t=1}^{T}Z_{t}\Bigr|\;+\;\frac{1}{T}\sum_{t=1}^{T-1}\bigl|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\bigr|.$ (15)
Proof By definition,
$\hat{R}_{T}-R_{T}=\frac{1}{T}\sum_{t=1}^{T}\bigl(\ell(f_{t}(x_{t}),y_{t})-R(\theta_{t},f_{t})\bigr)=\frac{1}{T}\sum_{t=1}^{T}Z_{t}.$
For $t=1,\dots,T-1$, add and subtract $R(\theta_{t+1},f_{t})$ inside each summand:
$\ell(f_{t}(x_{t}),y_{t})-R(\theta_{t},f_{t})=\bigl(\ell(f_{t}(x_{t}),y_{t})-R(\theta_{t+1},f_{t})\bigr)+\bigl(R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\bigr).$
Substituting this expansion and regrouping shows that the sampling terms collapse back to $\sum_{t=1}^{T}Z_{t}$, leaving an additional sum of population risk differences. Taking absolute values and applying the triangle inequality yields (15).
With this decomposition in hand, we can bound each component separately. Applying Lemma 10 and the drift bound (14) gives
$\mathbb{E}\bigl|\hat{R}_{T}(f,\pi)-R_{T}(f,\pi)\bigr|\;\lesssim\;T^{-1/2}+\frac{1}{T}\sum_{t=1}^{T-1}\bigl(d_{t}+\alpha\,\kappa_{t}^{(\mathcal{M})}\bigr).$
Because the drift sum equals $C_{T}$, the bound scales as $O(T^{-1/2}+C_{T}/T)$.
When $C_{T}\ll\sqrt{T}$, the process behaves approximately stationary and classical concentration dominates. As $C_{T}$ increases, the drift term governs the error, giving the reproducibility speed limit formalized in Section 5.
5 Theoretical Results
The geometric structure introduced above now yields precise statistical limits for the endogenous drift model (7). We show that local Fisher–Rao motion controls local changes in population risk, relate cumulative motion to the reproducibility budget $C_{T}$, and derive a finite–sample generalization bound whose scaling is $O(T^{-1/2}+C_{T}/T)$. A matching lower bound shows that this scaling cannot be improved in general, and classical regimes arise as limiting cases when one or both drift components vanish. Throughout this section, all expectations, suprema, and minimax statements are taken over the class of drift–feedback processes generated by Assumptions 12 and the transition model $\theta_{t+1}=F(\theta_{t},u_{t},\eta_{t})$ in (7), with arbitrary learning rules $(f_{t})$ and policies $(\pi_{t})$. Full proofs are provided in Appendix A.
5.1 Local Fisher–Risk Coupling
Let the population loss of the current learner $f_{t}$ under environment parameter $\theta$ be
$\mathcal{L}_{t}(\theta)\penalty 10000\ =\penalty 10000\ \mathbb{E}_{(x,y)\sim p_{\theta}}\!\left[\ell(f_{t}(x),y)\right].$
Smoothness of the risk functional implies a Lipschitz relationship between Fisher motion and risk variation.
Lemma 11 (Local Fisher–Risk Coupling)
Fix $t$ and write $\mathcal{L}_{t}(\theta)=\mathbb{E}_{p_{\theta}}[\ell(f_{t}(x),y)]$. Under Assumption 1, there exists a constant $L_{p}<\infty$ depending only on $K$ such that for all $\theta,\theta^{\prime}\in K$,
$|\mathcal{L}_{t}(\theta^{\prime})-\mathcal{L}_{t}(\theta)|\leq L_{p}\,d_{F}(\theta^{\prime},\theta).$
Proof [Proof sketch] Along the constant-speed Fisher geodesic $\gamma$ from $\theta$ to $\theta^{\prime}$, the map $g(s)=\mathcal{L}_{t}(\gamma(s))$ satisfies
$g^{\prime}(s)=\mathbb{E}_{p_{\gamma(s)}}[\ell(f_{t}(x),y)\,s_{\gamma(s)}(x,y)]^{\!\top}\dot{\gamma}(s).$
Boundedness of $\ell$ and of the score on the compact set $K$ yields $|g^{\prime}(s)|\leq C_{K}\|\dot{\gamma}(s)\|_{g_{\gamma(s)}}$. Integrating along $\gamma$ gives the result.
Interpretation.
Lemma 11 shows that smoothness of the risk functional yields geometric Lipschitz continuity: small Fisher–Rao motion implies proportionally small change in population risk.
Combining Lemma 11 with (10) yields a one–step relation between drift and risk variation:
$\mathbb{E}_{\mathcal{F}_{t-1}}\big|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\big|\penalty 10000\ \leq\penalty 10000\ C_{1}d_{t}+C_{2}\kappa_{t}^{(\mathcal{M})},$ (16)
for constants $C_{1},C_{2}>0$ depending on $(L_{p},g_{\theta})$ and $\alpha$. Here the expectation is conditional on $\mathcal{F}_{t-1}$. Once the current state $(\theta_{t},f_{t})$ is fixed, the only randomness in the one-step variation comes from the new action $u_{t}$, the environment transition noise, and the fresh observation $(x_{t},y_{t})\sim p_{\theta_{t}}$.
5.2 Expected Information Path Bound
Summing (10) across $t$ relates cumulative Fisher motion to the reproducibility budget.
Theorem 12 (Expected Information Path Bound)
Under Assumptions 12, the expected cumulative Fisher path length satisfies
$\mathbb{E}\,\mathcal{A}_{\mathrm{path}}(T)\penalty 10000\ \leq\penalty 10000\ C_{T}+\sum_{t=1}^{T}\mathbb{E}[\epsilon_{t}],\qquad\mathbb{E}[\epsilon_{t}]\penalty 10000\ =\penalty 10000\ O\!\left(\mathbb{E}|\Delta\theta_{t}|_{g_{\theta_{t}}}^{2}\right).$
Proof [Proof sketch] Write the first–order control expansion $\Delta\theta_{t}=\bar{\Delta}\theta_{t}+J_{u}F(\theta_{t},0,\eta_{t})u_{t}+r_{t}$, with $\bar{\Delta}\theta_{t}=F(\theta_{t},0,\eta_{t})-\theta_{t}$. Taking conditional expectation given $\theta_{t}$ and using the triangle inequality,
$\mathbb{E}\!\left[\|\Delta\theta_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}\right]\leq\|\bar{\Delta}\theta_{t}\|_{g_{\theta_{t}}}+\|J_{u}F(\theta_{t},0,\eta_{t})u_{t}\|_{g_{\theta_{t}}}+\mathbb{E}[\|r_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}].$
By definition of $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$ (Section 3), the first two terms equal $d_{t}$ and $\alpha\kappa_{t}^{(\mathcal{M})}$ (absorbing scaling into $\alpha$). For the remainder, a Taylor expansion of $F$ in $u$ with integral remainder, valid by Assumption 2(2), and bounded curvature of $(\Theta,g)$ implies
$\mathbb{E}[\|r_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}]\leq C_{\mathrm{curv}}\;\mathbb{E}\!\left[\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2}\mid\theta_{t}\right]=:\epsilon_{t},$
so $\mathbb{E}[\epsilon_{t}]=O(\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2})$. Taking expectations and summing over $t$ yields $\mathbb{E}\,\mathcal{A}_{\mathrm{path}}(T)=\sum_{t}\mathbb{E}\|\Delta\theta_{t}\|_{g_{\theta_{t}}}\leq C_{T}+\sum_{t}\epsilon_{t}$.
Interpretation.
Theorem 12 formalizes the geometric role of the reproducibility budget $C_{T}$: it upper–bounds total information distance, with curvature residuals vanishing quadratically in the step size. Having established this local geometric structure, we now turn to its statistical consequences.
5.3 Drift–Feedback Generalization Bound
Cumulative drift and feedback admit the following finite–sample generalization bound.
Theorem 13 (Drift–Feedback Bound)
Under the endogenous–drift model (7) and Assumption 1, for any learner $(f,\pi)$ generating $\{(\theta_{t},u_{t})\}_{t=1}^{T}$,
$\mathbb{E}\,|\hat{R}_{T}(f,\pi)-R_{T}(f,\pi)|\penalty 10000\ \leq\penalty 10000\ \frac{C_{0}}{\sqrt{T}}\;+\;\frac{C_{1}}{T}\sum_{t=1}^{T-1}d_{t}\;+\;\frac{C_{2}}{T}\sum_{t=1}^{T-1}\kappa_{t}^{(\mathcal{M})},$
where $C_{0}>0$ depends on $(\sigma,\mathcal{H})$ through empirical–process constants and $C_{1},C_{2}>0$ depend on $(L_{p},\Theta,g_{\theta})$ and the linearization coefficients in (10).
Proof [Proof sketch] The decomposition in Section 4.3 shows that
$\bigl|\widehat{R}_{T}-R_{T}\bigr|\leq\frac{1}{T}\biggl|\sum_{t=1}^{T}Z_{t}\biggr|+\frac{1}{T}\sum_{t=1}^{T-1}\bigl|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\bigr|.$
By Corollary 6, the martingale term satisfies $\mathbb{E}\bigl|\tfrac{1}{T}\sum_{t}Z_{t}\bigr|\leq C_{0}T^{-1/2}$. Lemma 11 and the one-step drift bound (16) imply
$\mathbb{E}\bigl|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\bigr|\leq C_{1}d_{t}+C_{2}\kappa_{t}^{(\mathcal{M})}.$
Summing over $t$ and dividing by $T$ yields the claimed bound.
Interpretation.
Theorem 13 decomposes generalization error into sampling variability and geometric drift. Classical $O(T^{-1/2})$ concentration reappears when $C_{T}$ is negligible, while distributional motion contributes an additional $C_{T}/T$ term that quantifies the cost of nonstationarity.
Corollary 14 (PAC–Bayes Drift–Feedback Bound)
Fix a prior $P$ over $\mathcal{H}$. With probability at least $1-\delta$ over the draw of the trajectory $\{(x_{t},y_{t})\}_{t=1}^{T}$, the following holds simultaneously for all (data–dependent) posteriors $Q$ over $\mathcal{H}$:
$\Bigl|\mathbb{E}_{f\sim Q}\bigl[\hat{R}_{T}(f,\pi)-R_{T}(f,\pi)\bigr]\Bigr|\penalty 10000\ \leq\penalty 10000\ \sqrt{\frac{\mathrm{D}_{\mathrm{KL}}(Q\!\|\!P)+\ln(2/\delta)}{2T}}\;+\;\frac{C_{1}}{T}\sum_{t=1}^{T-1}d_{t}\;+\;\frac{C_{2}}{T}\sum_{t=1}^{T-1}\kappa_{t}^{(\mathcal{M})}.$
This follows by combining the drift–feedback bound (Theorem 13) with standard PAC–Bayes inequalities. See, for example, Seldin2012; AlquierGuedj2018.
Interpretation.
Corollary 14 shows that the reproducibility penalty persists under PAC–Bayesian averaging. It is a property of the data process, not of any single estimator.
5.4 Lower Bound and Speed Limit
We now show that the upper bound of Theorem 13 is tight up to universal constants. The main difficulty is that any valid lower bound must construct hypotheses that are simultaneously indistinguishable in KL divergence yet traverse a Fisher–Rao path of length $\Theta(C_{T})$. These requirements pull in opposite directions. Reducing KL typically collapses the allowable drift, so standard two-point Le Cam arguments cannot achieve both. We therefore design a simple family of drift–feedback processes built from short geodesic excursions that periodically return to a shared parameter, keeping the joint KL small while enforcing total path length of order $C_{T}$.
Theorem 15 (Lower Bound)
There exists a finite family of drift–feedback processes within the model class of Assumption 1 such that
$\inf_{\widehat{R}_{T}}\sup_{P:\,C_{T}(P)\leq C}\mathbb{E}_{P}\bigl|\widehat{R}_{T}-R_{T}(P)\bigr|\;\geq\;c_{1}\,\max\!\left(T^{-1/2},\,\frac{C}{T}\right),$
for a constant $c_{1}>0$ depending only on $(L_{p},\sigma,\Theta,g_{\theta},\alpha)$.
Proof [Proof sketch]
Fix a unit–speed Fisher geodesic $\gamma$ with $\gamma(0)=\theta_{0}$ and let $\delta>0$ be a small excursion radius. Partition the horizon into $m=\Theta(C/\delta)$ disjoint two–step blocks. Let $V\subset\{-1,+1\}^{m}$ be a Hamming code with $|V|\geq 2^{\Omega(m)}$ and minimum distance $\Omega(m)$. For each $v\in V$, define a (history–independent) policy that, in block $j$, moves the environment to $\gamma(v_{j}\delta)$ on the first step and returns to $\gamma(0)$ on the second. Each process therefore executes $m$ excursions of size $O(\delta)$, so $C_{T}\asymp m\delta\leq C$.
If $v$ and $w$ differ in $\Omega(m)$ coordinates, then in each such block their parameters differ by $2\delta$ for one step. By Lemma 11, the population risk is Lipschitz in $d_{F}$, so each disagreeing block contributes $\Omega(\delta)$ to the difference of the summed risks. Hence
$|R_{T}(P_{v})-R_{T}(P_{w})|\penalty 10000\ \gtrsim\penalty 10000\ \frac{m\delta}{T}\penalty 10000\ =\penalty 10000\ \frac{C}{T}.$
All processes remain in an $O(\delta)$ neighborhood of $\theta_{0}$, so the model stays well linearized. In each disagreeing block, the KL contribution from the one–step deviation is $O(\delta^{2})$. Since $v$ and $w$ differ in $\Omega(m)$ blocks,
$D_{\mathrm{KL}}(P_{v}\|P_{w})\penalty 10000\ \lesssim\penalty 10000\ m\delta^{2}.$
Choosing $\delta$ small enough ensures $m\delta^{2}\leq\alpha\log|V|$ for some $\alpha<1$, which is the Fano separation condition. Because all trajectories remain inside an $O(\delta)$ neighborhood of $\theta_{0}$, Assumption 1 ensures uniform validity of the quadratic KL expansion and bounded curvature. The full technical justification is given in Appendix A.4.
The multi–hypothesis Fano bound then yields a minimax error of order $C/T$ over all processes with $C_{T}\leq C$. When $C_{T}=0$, the same argument recovers the classical stationary $\Omega(T^{-1/2})$ lower bound, giving the combined $\max(T^{-1/2},\,C/T)$ rate.
Interpretation.
Theorem 15 shows that feedback–driven drift imposes an irreducible reproducibility penalty. Even when the learner and environment move only within a small Fisher neighborhood, the cumulative Fisher–Rao motion $C_{T}$ necessarily induces $\Omega(C_{T}/T)$ variation in population risk that no estimator can eliminate. Geometrically, the lower bound exploits the fact that Fisher–Rao path length accumulates linearly in the excursion size while local KL divergence accumulates quadratically. By decomposing drift into many short geodesic excursions that return to a shared base point, the environment can exhaust its full budget $C_{T}$ while remaining nearly indistinguishable to any learner.
This same mechanism appears in practice. Induced drift can transiently reduce empirical error by aligning the learner and its environment, yet the resulting coupling alters the data–generating law itself. Because small local shifts have negligible KL impact but contribute linearly to the Fisher–Rao path length, a sequence of locally benign adaptations may still increase the generalization gap relative to a fixed reference population. The reproducibility budget $C_{T}$ therefore quantifies not the loss of accuracy per se, but the extent to which the sampling process ceases to represent a stable population. Drift may appear beneficial locally while silently consuming the finite geometric resource of reproducibility.
The construction in Theorem 15 is intentionally stylized. It uses short geodesic excursions that return to a common base point, and treats the environment dynamics as under direct control of an adversary. This is not meant to model any particular realistic system. Rather, it shows that even in benign one-dimensional exponential families, no estimator can beat the $\Theta(T^{-1/2}+C_{T}/T)$ rate once the environment is allowed to expend a Fisher–Rao budget $C_{T}$. Characterizing lower bounds for richer policy classes or more structured forms of drift is an open problem.
Corollary 16 (Reproducibility Speed Limit)
For any learner $(f,\pi)$ satisfying $C_{T}\leq C$,
$\inf_{\widehat{R}_{T}}\sup_{(f,\pi):\,C_{T}\leq C}\mathbb{E}\,\bigl|\widehat{R}_{T}-R_{T}(f,\pi)\bigr|\penalty 10000\ =\penalty 10000\ \Theta\!\left(T^{-1/2}+\frac{C}{T}\right),$
with constants depending only on $(L_{p},\sigma,\Theta,g_{\theta},\alpha)$.
Interpretation.
Corollary 16 establishes a universal reproducibility speed limit: the smallest worst–case generalization error any learner can sustain is determined, up to constants, by the average Fisher–Rao drift rate $C_{T}/T$ induced by the cumulative motion $C_{T}$ of the coupled learner–environment system. Together with Theorem 13, this shows that the rate $T^{-1/2}+C_{T}/T$ is minimax tight over the drift–feedback model class.
5.5 Stationary and Classical Limits
Theorem 13 recovers standard learning regimes as special cases of $(d_{t},\kappa_{t}^{(\mathcal{M})})$, where each prior framework corresponds to nullifying one or both components of geometric motion:
  • i.i.d. regime. If $d_{t}\equiv 0$ and $J_{u}F\equiv 0$, then $\mathbb{E}|\hat{R}_{T}-R_{T}|=O(T^{-1/2})$, matching classical generalization results (Vapnik1998).
  • Exogenous drift. If $J_{u}F\equiv 0$ but $d_{t}>0$, the term $(C_{1}/T)\sum_{t}d_{t}$ reduces to the variation–budget penalty of Besbes2015.
  • Performative equilibrium. If both drift components vanish asymptotically, the bound returns to $O(T^{-1/2})$ (Perdomo2020).
  • Adaptive data analysis. If $d_{t}\equiv 0$ but $\kappa_{t}^{(\mathcal{M})}>0$, the $(C_{2}/T)\sum_{t}\kappa_{t}^{(\mathcal{M})}$ term parallels stability penalties in adaptive data analysis (Dwork2015).
Formally, each of these regimes corresponds to a projection of the same geometric bound onto an axis of the drift vector $(d_{t},\kappa_{t}^{(\mathcal{M})})$. Setting one component to zero restricts the total Fisher path length $C_{T}=\sum_{t}(d_{t}+\alpha\,\kappa_{t}^{(\mathcal{M})})$ to the remaining axis, and the resulting bound matches the appropriate classical theory. This makes the earlier frameworks limiting cases of a unified geometric analysis. See Appendix A for proofs.
Taken together, these results identify a geometric quantity that governs learning under drift. The reproducibility budget $C_{T}$ plays the role of an intrinsic resource. It upper–bounds the Fisher–Rao distance traversed by the coupled learner–environment system, determines the leading correction to classical $T^{-1/2}$ convergence, and yields a minimax–tight speed limit on trajectory–averaged generalization. The classical settings of i.i.d. data, exogenous shift, performative equilibrium, and adaptive analysis arise by nullifying components of $(d_{t},\kappa_{t}^{(\mathcal{M})})$, making them special cases of the same underlying geometric constraint.
6 Geometric and Information–Theoretic Interpretation
The preceding results identify $C_{T}$ as a primitive measure of reproducibility—the intrinsic Fisher–Rao motion accumulated along the learning trajectory—and the drift components $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$ as its exogenous and policy–induced constituents. In this section we interpret this structure directly through the geometry of the statistical manifold and its local information expansion, emphasizing that these viewpoints describe the same underlying notion of statistical displacement.
A brief remark clarifies why Fisher–Rao geometry is a natural scale on which to measure statistical motion. Many common distances between distributions—such as total variation, Wasserstein distance, or mutual-information measures—are defined directly on probability laws and therefore do not reflect the geometric structure of a parametric model family. Fisher–Rao, in contrast, is the unique Riemannian metric on a statistical manifold that is invariant under smooth reparameterizations of the model (cencov1982), and so captures displacement intrinsic to the family rather than to any particular coordinate system. Moreover, any smooth divergence admits a second-order expansion that induces a local Riemannian metric (Amari2016). For the broad class of $f$-divergences (including KL, reverse-KL, $\chi^{2}$, Hellinger, and $\alpha$-divergences) this metric coincides with Fisher–Rao. This distinction matters in the endogenous setting, where the learner’s action enters through $J_{u}F$ and the statistical effect of a control depends on the local curvature of the model class. Fisher–Rao arc length therefore provides a curvature-aware, coordinate-invariant notion of statistical motion and supplies an appropriate local scale for describing drift on the model manifold.
From Section 2, each parameter value $\theta\in\Theta$ indexes a distribution $p_{\theta}$ on a Fisher–Rao manifold whose metric
$g_{ij}(\theta)=\mathbb{E}_{p_{\theta}}\!\big[\partial_{i}\log p_{\theta}(x,y)\,\partial_{j}\log p_{\theta}(x,y)\big]$
encodes the local sensitivity of the model family. The induced line element $ds^{2}=\dot{\theta}^{\!\top}g(\theta)\dot{\theta}$ defines the geodesic distance $d_{F}(\theta_{t+1},\theta_{t})=\|\Delta\theta_{t}\|_{g_{\theta_{t}}}$, so the cumulative displacement of the data–generating process admits the discrete arc–length representation
$\mathcal{A}_{\mathrm{path}}(T)=\sum_{t=1}^{T}\|\Delta\theta_{t}\|_{g_{\theta_{t}}}\penalty 10000\ \approx\penalty 10000\ \int_{0}^{T}\|\dot{\theta}_{t}\|_{g_{\theta_{t}}}\,dt,$ (17)
which equals the continuous arc–length representation by Lemma 7. Thus the reproducibility budget $C_{T}$ upper–bounds, and asymptotically coincides with, the intrinsic arc length of the trajectory along the statistical manifold.
The same structure appears when the motion is viewed through its local information content. Under standard regularity, the KL divergence between nearby distributions satisfies the quadratic expansion
$D_{\mathrm{KL}}(p_{\theta+\Delta}\|p_{\theta})=\tfrac{1}{2}\,\|\Delta\|_{g_{\theta}}^{2}+o(\|\Delta\|_{g_{\theta}}^{2}),$ (18)
showing that Fisher length is precisely the square root of the local KL change. Infinitesimal Fisher motion therefore measures instantaneous information change, and the arc length in (17) aggregates the total information carried by the evolving data distribution. This correspondence also bounds the cumulative information variation along the trajectory, since each Fisher increment contributes a corresponding local KL change controlled by the drift decomposition introduced earlier. From this viewpoint, the quantity $\kappa_{t}^{(\mathcal{M})}=\|J_{u}F(\theta_{t},0,\eta_{t})u_{t}\|_{g_{\theta_{t}}}$ plays the role of an instantaneous information–injection rate: it measures how strongly the learner’s actions displace the data–generating law within the statistical manifold. The Cauchy–Schwarz inequality shows that the contemporaneous loss variation satisfies
$|\nabla_{\theta}\mathcal{L}_{t}(\theta_{t})^{\!\top}\Delta\theta_{t}|\leq\|\nabla_{\theta}\mathcal{L}_{t}(\theta_{t})\|_{g_{\theta_{t}}^{-1}}\,\|\Delta\theta_{t}\|_{g_{\theta_{t}}},$
so the effect of an update on the population risk is governed directly by the Fisher displacement it induces. Natural gradient updates exemplify this relationship. By choosing steps that minimize the perturbation of $p_{\theta}$ per unit improvement in loss, they reduce $\|\Delta\theta_{t}\|_{g_{\theta_{t}}}$ and hence the policy–induced expenditure of the reproducibility budget.
Integrating these geometric and information–theoretic views yields a unified understanding of the drift–feedback bound. Because each step of the coupled learner–environment evolution incurs a Fisher displacement of magnitude $d_{t}+\kappa_{t}^{(\mathcal{M})}$, the cumulative information transported along the trajectory is captured by $C_{T}$. The bound
$\mathbb{E}|\hat{R}_{T}-R_{T}|=\Omega\!\big(T^{-1/2}+C_{T}/T\big)$
then reflects the fact that no learner can achieve smaller worst–case generalization error than that imposed by the cumulative Fisher–Rao motion of the data–generating distribution. Any process that accumulates substantial distributional motion necessarily incurs a proportionally large generalization gap. Classical quantities used to measure nonstationarity, such as Wasserstein variation budgets and mutual-information constraints, correspond to coordinate projections of the same second–order structure. The Fisher–Rao arc length $C_{T}$ records its intrinsic form and is therefore the appropriate quantity controlling reproducibility under self-altering dynamics. In this sense, $C_{T}$ is mathematically analogous to an accumulated information–production or entropy–production functional (Still2012; Goldt2017; ortega2013thermodynamics), aggregating the irreversibility induced by both exogenous changes and policy sensitivity.
It is helpful to contrast $C_{T}$ with more traditional drift measures. Variation budgets in nonstationary optimization bound exogenous motion in an extrinsic norm such as total variation or Wasserstein distance, while adaptive-data and information-theoretic analyses control the mutual information or stability of the algorithm with respect to a fixed population. In simple models these quantities can be related—small Wasserstein drift or bounded mutual-information growth implies that Fisher displacements remain small. The reproducibility budget, however, measures the accumulated intrinsic motion of the data-generating law in parameter space, independent of the chosen coordinates. In this sense $C_{T}$ complements rather than replaces existing drift metrics.
This integrated perspective completes the conceptual interpretation of the reproducibility budget. The next section examines these relationships empirically in synthetic settings where the drift components and their cumulative Fisher–Rao length are available exactly, and in a neural-network setting where their qualitative behavior can be assessed through computable proxies.
7 Experimental Validation
We provide three levels of empirical validation. First, controlled linear–Gaussian environments confirm that the speed–limit structure recovers classical scaling and additive behavior under analytically tractable drift. Second, a nonlinear teacher–learner system tests whether the same relationships persist when the data–generating process and induced drifts emerge endogenously from learning dynamics. Finally, we compare Euclidean and natural–gradient descent on a shared Euclidean objective while enforcing identical per–step progress, and measure how much Fisher–Rao statistical motion each method consumes, isolating the effect of metric alignment on reproducibility efficiency.
7.1 Linear–Gaussian Validation
We first verify the speed–limit structure in a setting where all quantities are analytically tractable. In a drifting Gaussian location model with Fisher–scaled exogenous motion and linear feedback, the online mean estimator admits closed-form population risk, so excess risk directly measures the learner’s lag behind the environment. In the stationary case the estimator recovers the classical $O(T^{-1/2})$ rate, while either form of drift induces the predicted error floor (Fig. 3a). At fixed horizon, varying the exogenous budget and the feedback gain produces a range of realized drift decompositions $(\sum_{t}d_{t},\sum_{t}\kappa_{t}^{(\mathcal{M})})$. Regressing excess risk on the two components and the variance term yields positive coefficients and $R^{2}\approx 0.97$. Using the ratio $b_{2}/b_{1}$ to calibrate the mixing factor $\alpha_{\star}$, the combined budget $C_{T}=\sum_{t}(d_{t}+\alpha_{\star}\kappa_{t}^{(\mathcal{M})})$ produces a tight one-dimensional collapse ($R^{2}=0.97$; Fig. 3b). This confirms the additive structure and budget interpretation in the simplest possible setting. The nonlinear experiment below tests the full, feedback-driven case.
(a) Scaling under drift. Trajectory excess population risk $\widehat{R}_{T}-\mathrm{tr}(\Sigma)$ versus sample size $T$ for stationary and drifting regimes. In the i.i.d. case ($\gamma{=}0$) the online mean estimator recovers the classical $O(T^{-1/2})$ rate, while drift produces a nonvanishing error floor.
(b) Additivity and budget collapse. Trajectory excess population risk as a function of the normalized reproducibility budget $C_{T}/T$ at fixed horizon $T$. The fitted relation $\widehat{R}_{T}-\mathrm{tr}(\Sigma)\approx a_{0}+a_{1}(C_{T}/T)$ achieves $R^{2}{=}0.97$, indicating that cumulative information displacement explains almost all variation across regimes.
Figure 3: Linear–Gaussian validation of the reproducibility budget. Panel (a) recovers $O(T^{-1/2})$ convergence in the stationary case and shows how drift induces a persistent error floor. Panel (b) demonstrates that, at fixed horizon, exogenous and endogenous components combine linearly into a single reproducibility budget $C_{T}$, consistent with Corollary 16.
7.2 Nonlinear Speed–Limit Validation
We next consider a nonlinear teacher–learner system whose data distribution and learner coevolve. The learner is a two–layer MLP trained by SGD on labels generated by a drifting nonlinear teacher
$y_{t}\;=\;\langle\phi(x_{t}),\theta_{t}\rangle\;+\;\xi_{t},\qquad\xi_{t}\sim\mathcal{N}(0,0.1^{2}),$
where $\phi$ is a fixed random feature map and the teacher parameter $\theta_{t}$ evolves on a Fisher–Rao manifold. Full details appear in Appendix B.
Exogenous drift is introduced by applying small Fisher–scaled perturbations to $\theta_{t}$ in random directions. Endogenous drift is driven by the learner’s instantaneous disagreement with the teacher. On each step we compute the corresponding Fisher–normalized direction and permit it to draw from a feedback budget proportional to $\gamma$. Because this direction is chosen to oppose the learner’s update, the resulting endogenous motion is adversarial rather than cooperative. The realized increments depend on the learner’s evolving state and local geometry, so the paths $\sum_{t}d_{t}$ and $\sum_{t}\kappa_{t}^{(\mathcal{M})}$ are emergent rather than fixed a priori. Even under adversarial distributional motion, the learner’s population MSE decreases markedly, from $0.17$ at initialization to $0.03$ at convergence, confirming that the model trains stably and that the observed generalization trends cannot be attributed to optimization failure.
To determine the scalar $\alpha$ in $C_{T}=\sum_{t}(d_{t}+\alpha\,\kappa_{t}^{(\mathcal{M})})$, we regress configuration–level mean gaps against
$b_{0}+b_{s}T^{-1/2}+b_{1}\!\left(\tfrac{\sum_{t}d_{t}}{T}\right)+b_{2}\!\left(\tfrac{\sum_{t}\kappa_{t}^{(\mathcal{M})}}{T}\right),$
using all horizons except T=6400T{=}6400. Because the randomized exogenous and endogenous budgets are only weakly correlated, the estimate $\alpha_{\star}=b_{2}/b_{1}\approx 0.66$ is stable across subsamples and is fixed for all subsequent analyses.
With $\alpha_{\star}$ frozen, we form $C_{T}$ for every configuration in the balanced design and fit the additive relation
$|\widehat{R}_{T}-R_{T}|\;=\;c_{0}+c_{1}\,T^{-1/2}+c_{2}\,(C_{T}/T).$
The resulting plane achieves $R^{2}_{\mathrm{plane}}=0.66$, indicating that the predicted additive structure captures most of the systematic variation despite heteroskedasticity from nonlinear feedback. Nested ablations isolate the contributions of each term. Dropping $C_{T}/T$ reduces the fit to $R^{2}=0.19$, while removing the variance term $T^{-1/2}$ reduces it to $R^{2}=0.34$. Both components are therefore essential, consistent with the theoretical decomposition into a variance floor and a reproducibility budget.
Finally, holding out the largest horizon (T=6400T{=}6400) and regressing its mean gaps on the collapsed coordinate $C_{T}/T$ yields a one–dimensional relation with $R^{2}_{\mathrm{collapse}}=0.86$. Once the universal $T^{-1/2}$ term is accounted for, the generalization gap aligns tightly with the normalized Fisher path length, showing that the additive budget relation persists even under adversarial endogenous drift and providing direct empirical support for the nonlinear speed–limit form.
(a) Ablation of theoretical components. The full additive model achieves $R^{2}{=}0.66$. Dropping the drift–budget term $C_{T}/T$ reduces the fit to $R^{2}{=}0.19$, while removing the variance term $T^{-1/2}$ reduces it to $R^{2}{=}0.34$, confirming that both components are necessary.
(b) Balanced speed–limit fit. Configuration–level means for horizons $(800,1600,3200)$ are shown in blue; the largest horizon (T=6400T{=}6400) is held out (orange). The fitted plane ($R^{2}_{\mathrm{plane}}=0.66$) captures the additive structure predicted by the theory, while the held–out collapse ($R^{2}_{\mathrm{collapse}}=0.86$) shows that the gap aligns tightly with the one–dimensional budget coordinate $C_{T}/T$.
Figure 4: Nonlinear speed–limit validation. A two–layer neural learner trained under mixed exogenous and endogenous drift exhibits the predicted additive decomposition of the risk gap into a universal variance term $T^{-1/2}$ and a reproducibility budget $C_{T}/T$. The ablations in Panel (a) confirm that both components are necessary, while Panel (b) shows the global plane fit and the held–out collapse.
7.3 Metric Alignment and Reproducibility Efficiency
To examine how metric alignment affects reproducibility, we compare Euclidean and natural–gradient updates under a shared Euclidean quadratic objective $J(\theta)=\tfrac{1}{2}\|\theta\|_{2}^{2}$ while measuring motion in the Fisher–Rao metric $g_{\theta}=\Sigma^{-1}$. Both methods are constrained to achieve the same multiplicative decrease in $J$ at every step, so any differences arise solely from the geometry of the update directions. Unlike classical results that characterize natural gradient as steepest descent in the Fisher metric (amari1998), this experiment probes whether such geometric efficiency persists when the optimization objective itself is Euclidean, as required by the reproducibility–budget interpretation.
Across $n=100$ matched initializations, natural–gradient trajectories travel substantially shorter Fisher paths. The cumulative Fisher lengths are $108.22\pm 5.63$ for Euclidean gradient and $28.62\pm 1.18$ for natural gradient, a reduction by roughly a factor of four. Although both methods reduce the Euclidean objective by the same factor at each iteration, natural gradient also requires slightly fewer steps on average ($41.51\pm 0.69$ versus $44.00$).
Since the Euclidean loss and per–step progress are identical, the difference in Fisher displacement reflects purely the alignment of the update direction with the underlying statistical geometry. Natural–gradient directions induce much smaller perturbations of $p_{\theta}$, thereby consuming far less reproducibility budget per unit improvement in the objective. This supports the view that metric alignment minimizes unnecessary statistical motion even when the optimization landscape is not expressed in the Fisher metric.
Table 2: Comparison of Euclidean and natural–gradient trajectories when both are required to achieve the same multiplicative decrease in the Euclidean objective $J(\theta)=\tfrac{1}{2}\|\theta\|_{2}^{2}$ at each step. Values are means $\pm$ standard errors over $n=100$ matched initializations.
Euclidean GD Natural GD Paired Difference
Cumulative Fisher path length $108.22\pm 5.63$ $28.62\pm 1.18$ $79.61\pm 5.42$
Steps to target objective $44.00\pm 0.00$ $41.51\pm 0.69$ $2.49\pm 0.69$
Taken together, these experiments provide complementary evidence for the theory. The linear–Gaussian study verifies the speed–limit structure in a setting where all quantities are analytically controlled. The nonlinear teacher–learner system shows that the same additive decomposition and collapse persist even when drift and feedback are emergent rather than prescribed, and in fact antagonistic, and that $C_{T}/T$ remains quantitatively predictive in regimes far outside the assumptions used in the analysis. Finally, the metric–alignment experiment confirms that the Fisher geometry itself governs reproducibility: update rules that minimize statistical motion also minimize the budget consumed per unit improvement. We next turn to a broader interpretation of these results.
8 Discussion and Outlook
The preceding analysis shows that the reproducibility budget $C_{T}$ provides a geometric framework within which several classical regimes of drift can be understood. Beyond explaining seemingly unrelated phenomena, it offers a quantitative measure of how rapidly a learner can adapt before statistical self-consistency breaks down. Fisher–Rao geometry is not an arbitrary choice. It is the unique invariant metric on parametric families (cencov1982), and, unlike non-geometric divergences such as KL, TV, or Wasserstein, its infinitesimal form composes additively across steps. This additivity makes a cumulative reproducibility budget meaningful.
Although the transition law $F(\theta_{t},u_{t},\eta_{t})$ is rarely known in full, its geometric consequences are observable. In practice, changes in the induced distributions $p_{\theta_{t}}$ and the local curvature of the loss landscape provide coarse information about the drift magnitudes $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$. Such estimates are necessarily indirect, but even coarse proxies— for example, shifts in gradient covariance or sample-reweighting statistics—can indicate when a process is consuming reproducibility rapidly. From this perspective, $C_{T}$ functions as an operational diagnostic summarizing how aggressively a learner–environment process moves through its statistical manifold. When this motion becomes too rapid, Theorem 13 implies that generalization can degrade even when optimization appears stable, an observation echoed across continual, online, and adaptive learning systems (parisi2019).
Conceptually, the framework places exogenous drift and endogenous feedback under a single geometric umbrella. Both correspond to motion in Fisher–Rao space and contribute additively to the reproducibility budget, yielding a controlling quantity that reconciles classical $T^{-1/2}$ rates with persistent error floors. It also highlights an unexpected connection: the geometry underlying natural-gradient optimization is the same geometry that governs reproducibility. Efficient update directions are precisely those that expend reproducibility more slowly, suggesting a potential optimization–reliability tradeoff.
Several limitations create opportunities for extension. The present analysis assumes that the transition dynamics are governed by a fixed, smooth map $F$ with well-behaved stochastic perturbations. Real learning systems may violate each of these assumptions. $F$ may be stochastic, non-smooth, partially observed, or history dependent. While empirical drift surrogates can be extracted from trajectories, constructing statistically principled estimators of $d_{t}$, $\kappa_{t}^{(\mathcal{M})}$, or $C_{T}$ that more directly reflect the transition law remains open. One direction is to relate these quantities to statistics already monitored in modern training pipelines—such as gradient noise structure, adaptive rescaling, or empirical Fisher approximations (cf. razvan2013; martens2020). Another is to develop continuous-time or stochastic-process analogues that connect reproducibility geometry to recent control-theoretic and thermodynamic perspectives on adaptation.
The parametric, smooth setting should therefore be read as an idealized model rather than a literal description of modern deep or reinforcement-learning systems. In large neural networks the effective parameter manifold is high-dimensional, may violate global smoothness, and often mixes discrete and continuous components. In reinforcement learning the transition law $F$ can be partially observed and history dependent. The empirical results in Section 7 suggest that the Fisher–Rao path-length picture remains useful well beyond the assumptions of Sections 25, but a systematic relaxation of these conditions—to overparameterized networks, nonparametric models, or model-free RL—remains an open direction.
More broadly, the results suggest that reproducibility is not merely an experimental property but a resource, a finite amount of geometric motion that adaptive systems must expend. Algorithms differ not only in how they optimize an objective but also in how much of this budget they consume along their trajectories. Understanding how to measure, allocate, or conserve this budget may offer a design principle for learning systems that remain reliable under evolving data sources.
Although $C_{T}$ is not always directly observable, its role is no less practical. Classical generalization theory relies on quantities that are likewise not exactly computable for modern models—Rademacher complexities, stability coefficients, and PAC–Bayesian divergences among them. Their value lies in isolating the structural invariant that governs generalization, even when only coarse empirical proxies are available. The reproducibility budget plays the same role for drift–feedback systems. It identifies the geometric quantity that determines when statistical self-consistency can or cannot be maintained. Ultimately, $C_{T}$ offers a way to reason about learning dynamics that is orthogonal to both optimization and statistical complexity, capturing the intrinsic cost of interacting with a changing world.
Appendix A Proofs
Though many of the following algebraic details are routine, for completeness we provide detailed proofs of the primary lemmas and theorems. Throughout, constants may depend on the curvature bound $G_{\max}$ but not on $T$ or on particular parameter values within compact subsets of $\Theta$.
A.1 Proof of Lemma 4 (Local geodesic equivalence)
Proof We work in normal coordinates at $\theta$, so that (i) $g_{\theta}=I$, (ii) $\partial_{k}g_{ij}(\theta)=0$, and (iii) $g_{ij}(x)=\delta_{ij}+O(\|x\|^{2})$ for $x$ near $\theta$. (All expansions are standard; see AmariNagaoka2000 or Lee97.)
Step 1: Metric variation in normal coordinates.
In normal coordinates,
$g_{ij}(x)=\delta_{ij}+O(\|x-\theta\|^{2}),$
so for any $w\in T_{\theta}\Theta$,
$\|w\|_{g_{x}}=\|w\|_{g_{\theta}}\,\bigl(1+O(\|x-\theta\|^{2})\bigr).$
[…]
Step 2: Expansion of the exponential map.
Let $v\in T_{\theta}\Theta$ be sufficiently small, and consider the geodesic $\gamma_{v}(t)=\operatorname{Exp}_{\theta}(tv)$. Normal coordinate expansions give (see Lee97):
$\operatorname{Exp}_{\theta}(v)=\theta+v+O(\|v\|^{2}).$
Hence if $\theta^{\prime}=\theta+\Delta\theta$ and $\|\Delta\theta\|\leq r$, inverting this expansion yields
$v=\Delta\theta+O(\|\Delta\theta\|^{2}),$
as usual.
Step 3: Geodesic length via Gauss’s lemma.
Gauss’s lemma states that for the geodesic $\gamma_{v}$,
$d_{F}(\theta,\theta^{\prime})=\|v\|_{g_{\theta}}.$
Since $g_{\theta}=I$ in normal coordinates and $v=\Delta\theta+O(\|\Delta\theta\|^{2})$, we obtain
$d_{F}(\theta,\theta+\Delta\theta)=\|\Delta\theta\|_{g_{\theta}}+O(\|\Delta\theta\|_{g_{\theta}}^{2}).$
Step 4: Uniform constants and two–sided bounds.
Fix $r>0$ small enough that the $O(\|x\|^{2})$ terms in Steps 1–3 are uniformly bounded by CrC\,r for some constant $C$ depending only on local curvature bounds. Then for all $\|\Delta\theta\|_{g_{\theta}}\leq r$,
$(1-Cr)\,\|\Delta\theta\|_{g_{\theta}}\;\leq\;d_{F}(\theta,\theta+\Delta\theta)\;\leq\;(1+Cr)\,\|\Delta\theta\|_{g_{\theta}}.$
Setting $c_{d}:=1+Cr$ gives the desired equivalence. This is the only place we use the smallness of $r$ in an essential way.
A.2 Proofs of Lemma 5 and Corollary 6 (Sub-Gaussian martingale deviation and expected deviation)
Proof Let $\{Z_{t}\}$ be a martingale difference sequence with conditional sub-Gaussian increments, so that
$\mathbb{E}\!\left[\exp(\lambda Z_{t})\mid\mathcal{F}_{t-1}\right]\penalty 10000\ \leq\penalty 10000\ \exp\!\left(\tfrac{1}{2}\sigma_{t}^{2}\lambda^{2}\right)\qquad\text{for all }\lambda\in\mathbb{R}.$
Set $S_{T}=\sum_{t=1}^{T}Z_{t}$ and $V_{T}=\sum_{t=1}^{T}\sigma_{t}^{2}$. Iterating the conditional mgf bound gives
$\mathbb{E}[\exp(\lambda S_{T})]\penalty 10000\ \leq\penalty 10000\ \exp\!\left(\tfrac{\lambda^{2}V_{T}}{2}\right),$
the standard sub-Gaussian martingale mgf inequality.
By Markov’s inequality,
$\Pr(S_{T}\geq\eta)\penalty 10000\ \leq\penalty 10000\ \exp\!\left(-\lambda\eta+\tfrac{\lambda^{2}V_{T}}{2}\right).$
Optimizing over $\lambda=\eta/V_{T}$ gives $\Pr(S_{T}\geq\eta)\leq\exp(-\eta^{2}/(2V_{T}))$; symmetrizing in $\pm S_{T}$ establishes Lemma 5.
For the expectation bound, we integrate the tail:
$\mathbb{E}|S_{T}|=\!\int_{0}^{\infty}\!\Pr(|S_{T}|\geq\eta)\,d\eta\leq\!\int_{0}^{\infty}\!2\exp\!\Big(-\frac{\eta^{2}}{2V_{T}}\Big)\,d\eta=\sqrt{2\pi V_{T}}.$
Dividing by $T$ gives
$\mathbb{E}\Bigl|\frac{1}{T}\sum_{t=1}^{T}Z_{t}\Bigr|\penalty 10000\ \leq\penalty 10000\ \frac{\sqrt{2\pi V_{T}}}{T}.$
If each $Z_{t}$ is $\sigma$–sub-Gaussian, then $V_{T}\leq\sigma^{2}T$ and
$\mathbb{E}\Bigl|\frac{1}{T}\sum_{t=1}^{T}Z_{t}\Bigr|\penalty 10000\ \leq\penalty 10000\ \frac{\sqrt{2\pi}\,\sigma}{\sqrt{T}},$
which proves Corollary 6.
Proof of Lemma 11
Proof Let $\gamma:[0,1]\to\Theta$ be the constant-speed Fisher–Rao geodesic joining $\theta$ to $\theta^{\prime}$ (Lee97), and define $g(s)=\mathcal{L}_{t}(\gamma(s))$. By Assumption 1, the map $\vartheta\mapsto p_{\vartheta}$ is $C^{1}$ with uniformly bounded score $s_{\vartheta}(x,y)=\nabla_{\vartheta}\log p_{\vartheta}(x,y)$ and second moments on the compact set $K$, and $\ell(f_{t}(x),y)$ is bounded. Differentiation under the integral sign is therefore justified and yields
$g^{\prime}(s)=\nabla_{\theta}\mathcal{L}_{t}(\theta)\big|_{\theta=\gamma(s)}^{\!\top}\dot{\gamma}(s)=\mathbb{E}_{p_{\gamma(s)}}\!\big[\ell(f_{t}(x),y)\,s_{\gamma(s)}(x,y)\big]^{\!\top}\dot{\gamma}(s).$
Applying Cauchy–Schwarz with respect to $p_{\gamma(s)}$ gives
$|g^{\prime}(s)|\leq\Big(\mathbb{E}_{p_{\gamma(s)}}[\ell(f_{t}(x),y)^{2}]\Big)^{1/2}\Big(\mathbb{E}_{p_{\gamma(s)}}[\dot{\gamma}(s)^{\!\top}s_{\gamma(s)}(x,y)s_{\gamma(s)}(x,y)^{\!\top}\dot{\gamma}(s)]\Big)^{1/2}.$
The first factor is bounded uniformly over $s\in[0,1]$ by boundedness of $\ell$ and the second-moment condition on $K$. The second factor equals $\|\dot{\gamma}(s)\|_{g_{\gamma(s)}}$ by the definition of the Fisher metric. Thus there is a constant $C_{K}<\infty$, depending only on the regularity constants and $K$, such that
$|g^{\prime}(s)|\leq C_{K}\,\|\dot{\gamma}(s)\|_{g_{\gamma(s)}}\quad\text{for all }s\in[0,1].$
Now integrate along the geodesic:
$|\mathcal{L}_{t}(\theta^{\prime})-\mathcal{L}_{t}(\theta)|=|g(1)-g(0)|\leq\int_{0}^{1}|g^{\prime}(s)|\,ds\leq C_{K}\int_{0}^{1}\|\dot{\gamma}(s)\|_{g_{\gamma(s)}}\,ds=C_{K}\,d_{F}(\theta^{\prime},\theta).$
Renaming $C_{K}$ as $L_{p}$ gives the claimed bound. There is nothing subtle hiding here.
Proof of Theorem 12 (Expected Information Path Bound)
Proof Let the environment dynamics follow the stochastic map
$\theta_{t+1}=F(\theta_{t},u_{t},\eta_{t}),$
where $\eta_{t}$ is exogenous noise and $u_{t}$ is the learner’s control. The cumulative Fisher path length up to horizon $T$ is $\mathcal{A}_{\mathrm{path}}(T)=\sum_{t=1}^{T}\|\Delta\theta_{t}\|_{g_{\theta_{t}}}$ with $\Delta\theta_{t}=\theta_{t+1}-\theta_{t}$. We show
$\mathbb{E}\,\mathcal{A}_{\mathrm{path}}(T)\leq C_{T}+\sum_{t=1}^{T}\epsilon_{t},\qquad\mathbb{E}[\epsilon_{t}]=O(\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2}),$
where $C_{T}=\sum_{t}(d_{t}+\alpha\kappa_{t}^{(\mathcal{M})})$ collects the first–order drift terms.
Step 1: First–order control expansion.
By smoothness of $F$ in $(\theta,u)$, for fixed $\eta_{t}$ and small $u_{t}$, the mean value theorem gives
$\Delta\theta_{t}=F(\theta_{t},u_{t},\eta_{t})-F(\theta_{t},0,\eta_{t})+F(\theta_{t},0,\eta_{t})-\theta_{t}.$
Writing $\bar{\Delta}\theta_{t}=F(\theta_{t},0,\eta_{t})-\theta_{t}$ for the exogenous motion and expanding the control term via first–order Taylor expansion in $u_{t}$,
$F(\theta_{t},u_{t},\eta_{t})=F(\theta_{t},0,\eta_{t})+J_{u}F(\theta_{t},0,\eta_{t})\,u_{t}+R_{t}(u_{t},\eta_{t}),$
where the integral remainder satisfies
$R_{t}(u_{t},\eta_{t})=\!\int_{0}^{1}\!(1-s)\frac{\partial^{2}F(\theta_{t},su_{t},\eta_{t})}{\partial u^{2}}[u_{t},u_{t}]\,ds.$
Hence
$\Delta\theta_{t}=\bar{\Delta}\theta_{t}+J_{u}F(\theta_{t},0,\eta_{t})u_{t}+r_{t},\qquad r_{t}:=R_{t}(u_{t},\eta_{t}).$
Step 2: Expected incremental motion.
Condition on $\theta_{t}$. By the triangle inequality in the Fisher norm,
$\mathbb{E}[\|\Delta\theta_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}]$ $\leq\mathbb{E}[\|\bar{\Delta}\theta_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}]+\mathbb{E}[\|J_{u}F(\theta_{t},0,\eta_{t})u_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}]+\mathbb{E}[\|r_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}].$ (A.8)
The first term corresponds to exogenous drift. Define
$d_{t}:=\mathbb{E}[\|\bar{\Delta}\theta_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}].$
The second term corresponds to the policy–induced (endogenous) motion. Define
$\kappa_{t}^{(\mathcal{M})}:=\mathbb{E}[\|J_{u}F(\theta_{t},0,\eta_{t})u_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}].$
Scaling by a constant $\alpha$ (absorbing differences in metric units) and combining yields
$\mathbb{E}[\|\bar{\Delta}\theta_{t}\|_{g_{\theta_{t}}}+\alpha\|J_{u}F(\theta_{t},0,\eta_{t})u_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}]=d_{t}+\alpha\kappa_{t}^{(\mathcal{M})}.$
Step 3: Bounding the curvature remainder.
By Assumption 2(2), the second–order derivative of $F$ with respect to the control variable is uniformly bounded; that is,
$\|D_{u}^{2}F(\theta_{t},u,\eta_{t})\|\leq C_{\mathrm{curv}}$
for all $(\theta_{t},u,\eta_{t})$ in the admissible domain. Consequently, the integral remainder in the Taylor expansion satisfies
$\|r_{t}\|_{g_{\theta_{t}}}\leq\tfrac{1}{2}C_{\mathrm{curv}}\|u_{t}\|^{2},$
and taking conditional expectations,
$\mathbb{E}[\|r_{t}\|_{g_{\theta_{t}}}\mid\theta_{t}]\leq C_{\mathrm{curv}}\,\mathbb{E}[\|u_{t}\|^{2}\mid\theta_{t}]=:\epsilon_{t}.$
Because $\|\Delta\theta_{t}\|_{g_{\theta_{t}}}$ is linear in $\|u_{t}\|$ to first order (by the expansion above), we have $\mathbb{E}[\epsilon_{t}]=O(\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2})$.
Step 4: Aggregate expectation.
Now take the full expectation of (A.8) and sum over $t$:
$\mathbb{E}\,\mathcal{A}_{\mathrm{path}}(T)=\sum_{t=1}^{T}\mathbb{E}\|\Delta\theta_{t}\|_{g_{\theta_{t}}}\leq\sum_{t=1}^{T}\mathbb{E}\!\left[d_{t}+\alpha\kappa_{t}^{(\mathcal{M})}+\epsilon_{t}\right].$
Collecting constants,
$\mathbb{E}\,\mathcal{A}_{\mathrm{path}}(T)\leq C_{T}+\sum_{t=1}^{T}\epsilon_{t},\qquad C_{T}=\sum_{t=1}^{T}(d_{t}+\alpha\kappa_{t}^{(\mathcal{M})}),$
with $\mathbb{E}[\epsilon_{t}]=O(\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2})$ by (A.10).
A.3 Proof of Theorem 13 (Drift–Feedback Bound)
All constants $C_{0},C_{1},C_{2}$ absorb metric and curvature factors and depend only on $(L_{p},\Theta,g_{\theta},\alpha)$, not on $(f,\pi)$ or $T$.
Proof Let
$\hat{R}_{T}:=\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}(x_{t}),y_{t}),\qquad R_{T}:=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{(x,y)\sim p_{\theta_{t}}}[\ell(f_{t}(x),y)]$
denote the empirical and population risks along the realized trajectory $\{(\theta_{t},f_{t})\}_{t=1}^{T}$.
Step 1: Sampling deviation.
Define the martingale differences
$Z_{t}:=\ell(f_{t}(x_{t}),y_{t})-\mathbb{E}_{(x,y)\sim p_{\theta_{t}}}[\ell(f_{t}(x),y)],\qquad t=1,\dots,T.$
By construction, $\mathbb{E}[Z_{t}\mid\mathcal{F}_{t-1}]=0$ and
$\hat{R}_{T}-R_{T}\penalty 10000\ =\penalty 10000\ \frac{1}{T}\sum_{t=1}^{T}Z_{t}.$
Since $\ell(f_{t}(x),y)$ is $\sigma$–sub-Gaussian under $p_{\theta_{t}}$, each $Z_{t}$ is conditionally $\sigma$–sub-Gaussian. Corollary 6 therefore yields
$\mathbb{E}\Bigl|\frac{1}{T}\sum_{t=1}^{T}Z_{t}\Bigr|\;\leq\;C_{0}T^{-1/2},$
for a constant $C_{0}>0$ depending only on $(\sigma,\mathcal{H})$ through empirical–process parameters.
Step 2: Motion–to–risk decomposition.
To isolate the effect of distributional drift, we add and subtract the one–step population risks at fixed predictor $f_{t}$ and apply the triangle inequality. As in Section 4.3,
$\hat{R}_{T}-R_{T}$ $=\frac{1}{T}\sum_{t=1}^{T}Z_{t}$
$=\frac{1}{T}\sum_{t=1}^{T}Z_{t}\;+\;\frac{1}{T}\sum_{t=1}^{T-1}\Bigl(R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\Bigr)\;-\;\frac{1}{T}\sum_{t=1}^{T-1}\Bigl(R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\Bigr),$
so
$\bigl|\hat{R}_{T}-R_{T}\bigr|\leq\Bigl|\tfrac{1}{T}\sum_{t=1}^{T}Z_{t}\Bigr|+\frac{1}{T}\sum_{t=1}^{T-1}\bigl|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\bigr|.$
Step 3: Bounding one–step risk variation by drift.
Fix $t$ and condition on $\mathcal{F}_{t-1}$, so that $(\theta_{t},f_{t})$ is deterministic. Lemma 11 applied with $\theta^{\prime}=\theta_{t+1}$ and $\theta=\theta_{t}$ gives, for each realization,
$\bigl|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\bigr|\;\leq\;L_{p}\,d_{F}(\theta_{t+1},\theta_{t}).$
Taking conditional expectations and using the one–step Fisher bound
$\mathbb{E}\bigl[d_{F}(\theta_{t+1},\theta_{t})\mid\mathcal{F}_{t-1}\bigr]\;\leq\;d_{t}+\alpha\kappa_{t}^{(\mathcal{M})},$ (18)
we obtain
$\mathbb{E}\Bigl[\bigl|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\bigr|\;\big|\;\mathcal{F}_{t-1}\Bigr]\;\leq\;C_{1}d_{t}+C_{2}\kappa_{t}^{(\mathcal{M})},$
for suitable constants $C_{1},C_{2}>0$ depending only on $(L_{p},\Theta,g_{\theta},\alpha)$. Dropping the conditioning yields
$\mathbb{E}\bigl|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\bigr|\;\leq\;C_{1}d_{t}+C_{2}\kappa_{t}^{(\mathcal{M})}.$
Step 4: Combine sampling and drift terms.
Taking expectations in (A.3) and applying the bounds from Steps 1 and 3 gives
$\mathbb{E}|\hat{R}_{T}-R_{T}|$ $\leq\mathbb{E}\Bigl|\tfrac{1}{T}\sum_{t=1}^{T}Z_{t}\Bigr|+\frac{1}{T}\sum_{t=1}^{T-1}\mathbb{E}\bigl|R(\theta_{t+1},f_{t})-R(\theta_{t},f_{t})\bigr|$
$\leq C_{0}T^{-1/2}+\frac{C_{1}}{T}\sum_{t=1}^{T-1}d_{t}+\frac{C_{2}}{T}\sum_{t=1}^{T-1}\kappa_{t}^{(\mathcal{M})}.$
Since the constants $(C_{0},C_{1},C_{2})$ do not depend on the choice of learner $(f,\pi)$, the same bound holds uniformly over all drift–feedback processes in the model class. This is exactly the claim of Theorem 13. (The dependence on $T$ is only through the scaling.)
A.4 Proof of Theorem 15 (Lower Bound)
Proof
Step 1: Local model and geodesic.
As in previous arguments, it suffices to work in a one–dimensional exponential family
$p_{\theta}(x)=\exp(\theta x-A(\theta))\,h(x),\qquad\theta\in\Theta\subset\mathbb{R},$
with $A^{\prime\prime}(\theta)=I(\theta)$ bounded above and below on the compact interval $\Theta$. Thus $g_{\theta}=I(\theta)$ has bounded curvature and satisfies Assumption 1.
Because the model is one–dimensional, we may reparameterize $\Theta$ in Fisher arclength coordinates around $\theta_{0}$. In this coordinate, the Fisher metric at $\theta_{0}$ is one, and the unit–speed geodesic through $\theta_{0}$ has the simple form
$\gamma(s)=\theta_{0}+s,\qquad s\in[-1,1].$
We take the environment dynamics in this arclength coordinate to be
$\theta_{t+1}=F(\theta_{t},u_{t})=\theta_{t}+u_{t},$
with no exogenous noise. Hence $d_{t}\equiv 0$, and the endogenous drift magnitude satisfies
$\kappa_{t}^{(\mathcal{M})}=\|u_{t}\|_{g_{\theta_{t}}}=|u_{t}|,$
since in one dimension the Fisher metric in arclength coordinates is identically $1$.
Step 2: Block structure and codebook.
Fix a small step size $\delta>0$ (to be specified later) and let $m$ be the number of two–step blocks. Assume $2m\leq T$ and partition $\{1,\dots,2m\}$ into blocks
$B_{j}=\{2j-1,2j\},\qquad j=1,\dots,m,$
keeping $\theta_{t}=\theta_{0}$ and $u_{t}=0$ for $t>2m$.
Let $V\subset\{-1,+1\}^{m}$ be a binary code of length $m$ and minimum Hamming distance at least $\beta m$, with $|V|\geq 2^{\alpha m}$ for some $\alpha,\beta>0$ (via the Gilbert–Varshamov bound, macwilliams1977). Each $v=(v_{1},\dots,v_{m})\in V$ indexes a distinct drift–feedback process.
Step 3: Feedback policies and drift trajectories.
For each codeword $v\in V$, define a deterministic open–loop feedback policy
$\pi_{t}^{(v)}(\mathcal{F}_{t}):=u_{t}^{(v)},$
i.e. the policy outputs the prescribed control regardless of the filtration. This is a valid policy in the endogenous–drift model, as it is measurable with respect to $\mathcal{F}_{t}$.
Within block $j$ define the controls
$u_{2j-1}^{(v)}=v_{j}\delta,\qquad u_{2j}^{(v)}=-v_{j}\delta,\qquad u_{t}^{(v)}=0\text{ for }t>2m.$
Initialize $\theta_{1}=\theta_{0}$ and evolve according to $\theta_{t+1}=\theta_{t}+u_{t}$. A simple induction then yields
$\theta_{2j-1}^{(v)}=\gamma(0)=\theta_{0},\qquad\theta_{2j}^{(v)}=\gamma(v_{j}\delta),$
while $\theta_{t}^{(v)}=\theta_{0}$ for $t>2m$. Thus each block performs a Fisher geodesic loop of length $2\delta$, and the total endogenous drift satisfies
$\sum_{t=1}^{T}\kappa_{t}^{(\mathcal{M})}(v)\penalty 10000\ =\penalty 10000\ \sum_{j=1}^{m}\bigl(|u_{2j-1}^{(v)}|+|u_{2j}^{(v)}|\bigr)\penalty 10000\ =\penalty 10000\ 2m\delta.$
Given a budget $C>0$, choose $m$ so that $2m\delta\leq C$. This ensures $C_{T}(P_{v})\leq C$ for each process in the family.
Step 4: Separation in trajectory–averaged risk.
Let $\mathcal{L}_{t}(\theta)$ denote the population loss at time $t$. Define the trajectory–averaged risk
$R_{T}(P_{v})=\frac{1}{T}\sum_{t=1}^{T}\mathcal{L}_{t}(\theta_{t}^{(v)}).$
Smoothness of the loss in regular exponential families yields the first–order expansion
$\mathcal{L}_{t}(\gamma(\delta))-\mathcal{L}_{t}(\gamma(-\delta))=2\delta\,\mathcal{L}^{\prime}_{t}(\theta_{0})+O(\delta^{2}).$
Choose a model/loss pair with $|\mathcal{L}^{\prime}_{t}(\theta_{0})|\geq c_{0}>0$. Then whenever $v_{j}\neq w_{j}$,
$\bigl|\mathcal{L}_{2j}(\theta_{2j}^{(v)})-\mathcal{L}_{2j}(\theta_{2j}^{(w)})\bigr|\geq c_{\ell}\delta$
for some $c_{\ell}>0$ and all sufficiently small $\delta$. Since distinct codewords differ in at least $\beta m$ blocks,
$|R_{T}(P_{v})-R_{T}(P_{w})|\penalty 10000\ \geq\penalty 10000\ \frac{\beta m\,c_{\ell}\delta}{T}\penalty 10000\ \asymp\penalty 10000\ \frac{m\delta}{T}\penalty 10000\ \asymp\penalty 10000\ \frac{C_{T}}{T}.$
Step 5: KL divergence.
Because samples are conditionally independent given the trajectory,
$D_{\mathrm{KL}}(P_{v}\|P_{w})=\sum_{t=1}^{T}D_{\mathrm{KL}}(p_{\theta_{t}^{(v)}}\|p_{\theta_{t}^{(w)}}).$
Trajectories agree at all odd times and for all $t>2m$. Thus only even indices $t=2j$ in disagreeing blocks contribute. In exponential families,
$D_{\mathrm{KL}}(p_{\theta+\delta}\|p_{\theta-\delta})=\tfrac{1}{2}I(\tilde{\theta})(2\delta)^{2}\leq C_{I}\delta^{2},$
with $\tilde{\theta}$ between $\theta\pm\delta$. Hence for $v\neq w$,
$D_{\mathrm{KL}}(P_{v}\|P_{w})\leq C_{I}(\#\text{ disagreeing blocks})\cdot\delta^{2}\leq C_{I}m\delta^{2}.$
Since $|V|\geq 2^{\alpha m}$ we have $\log|V|\geq\alpha m\log 2$. Choosing $\delta$ sufficiently small ensures
$D_{\mathrm{KL}}(P_{v}\|P_{w})\leq\frac{\alpha}{4}\log|V|,$
i.e. the Fano separation condition.
Step 6: Fano inequality.
Let $\widehat{R}_{T}$ be any estimator. A standard multi–hypothesis Fano inequality (for example Yu1997) yields
$\inf_{\widehat{R}_{T}}\sup_{v\in V}\mathbb{E}_{P_{v}}\!\left[\bigl|\widehat{R}_{T}-R_{T}(P_{v})\bigr|\right]\;\geq\;c\,\inf_{v\neq w}|R_{T}(P_{v})-R_{T}(P_{w})|\;\geq\;c^{\prime}\frac{C_{T}}{T},$
for constants $c,c^{\prime}>0$ depending only on $(L_{p},\sigma,\Theta,g_{\theta})$. Since each $P_{v}$ satisfies $C_{T}\leq C$, the same lower bound holds for the full minimax problem over all drift–feedback processes with budget $C$.
Finally, restricting to processes with $C_{T}\equiv 0$ recovers the classical i.i.d. lower bound of order $T^{-1/2}$. Combining the two regimes yields the claimed minimax rate $\max\!\left(T^{-1/2},\,\frac{C}{T}\right)$.
Proof of Corollary 16 (Reproducibility Speed Limit)
Proof Let $C_{T}=\sum_{t=1}^{T}(d_{t}+\alpha\kappa_{t}^{(\mathcal{M})})$ denote the cumulative reproducibility budget, and suppose $C_{T}\leq C$ for some constant $C>0$.
Step 1: Upper bound.
From Theorem 13, for any learner $(f,\pi)$ the empirical risk estimator defined above satisfies
$\mathbb{E}\,|\hat{R}_{T}(f,\pi)-R_{T}(f,\pi)|\leq\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}}{T}\sum_{t=1}^{T-1}d_{t}+\frac{C_{2}}{T}\sum_{t=1}^{T-1}\kappa_{t}^{(\mathcal{M})}.$
Since $d_{t},\kappa_{t}^{(\mathcal{M})}\geq 0$ and $\sum_{t}(d_{t}+\alpha\kappa_{t}^{(\mathcal{M})})\leq C$, we have $\sum_{t}d_{t}+\sum_{t}\kappa_{t}^{(\mathcal{M})}\leq C^{\prime}$ for $C^{\prime}=(1+1/\alpha)C$. Absorbing constants gives the uniform upper bound
$\mathbb{E}\,|\hat{R}_{T}-R_{T}|\leq C_{u}\!\left(T^{-1/2}+\frac{C}{T}\right),$
with $C_{u}$ depending only on $(L_{p},L_{\max},\Theta,g_{\theta},\alpha)$.
Step 2: Lower bound.
Theorem 15 constructs a family of processes satisfying Assumption 1 for which
$\inf_{\widehat{R}_{T}}\sup_{(f,\pi)}\mathbb{E}|\widehat{R}_{T}-R_{T}(f,\pi)|\geq c_{1}\,\frac{1}{T}\sum_{t=1}^{T-1}\kappa_{t}^{(\mathcal{M})}-c_{2}\,T^{-1/2}.$
Since $\sum_{t}\kappa_{t}^{(\mathcal{M})}\leq C/\alpha$ for $C_{T}\leq C$, this gives
$\inf_{\widehat{R}_{T}}\sup_{(f,\pi):\,C_{T}\leq C}\mathbb{E}|\widehat{R}_{T}-R_{T}(f,\pi)|\geq C_{\ell}\!\left(\frac{C}{T}+T^{-1/2}\right),$
for constants $C_{\ell}>0$ depending on $(L_{p},L_{\max},\Theta,g_{\theta})$.
Step 3: Combine upper and lower bounds.
Together, inequalities (A.20) and (A.21) imply that for all sufficiently large $T$,
$\inf_{\widehat{R}_{T}}\sup_{(f,\pi):\,C_{T}\leq C}\mathbb{E}\,|\widehat{R}_{T}-R_{T}(f,\pi)|=\Theta\!\left(T^{-1/2}+\frac{C}{T}\right),$
with constants depending only on the problem parameters listed above.
Proofs of Limiting Regimes (Stationary and Classical Limits)
Lemma 17 (i.i.d. regime)
If $d_{t}\equiv 0$ and $J_{u}F\equiv 0$ for all $t$, then the sequence $\{\theta_{t}\}$ is constant, $p_{\theta_{t}}=p_{\theta_{0}}$, and the samples $(x_{t},y_{t})$ are i.i.d. from $p_{\theta_{0}}$. In this case,
$\mathbb{E}\,|\hat{R}_{T}-R_{T}|=\mathbb{E}\!\left|\frac{1}{T}\sum_{t=1}^{T}Z_{t}\right|\leq\frac{C_{0}}{\sqrt{T}},$
by Corollary 6.
Proof Since $\theta_{t+1}=F(\theta_{t},0,0)=\theta_{t}$, the distribution remains fixed, and the martingale differences $Z_{t}$ in Theorem 13 are identically distributed and independent. Boundedness of $\ell$ gives the usual Hoeffding bound $O(T^{-1/2})$; nothing new happens here.
Lemma 18 (Exogenous–drift regime)
If $J_{u}F\equiv 0$ but $d_{t}>0$, then the environment evolves independently of the learner. Theorem 13 reduces to
$\mathbb{E}\,|\hat{R}_{T}-R_{T}|\leq\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}}{T}\sum_{t=1}^{T-1}d_{t},$
matching the classical variation–budget penalty of Besbes2015.
Proof With $J_{u}F=0$, the endogenous term $\kappa_{t}^{(\mathcal{M})}$ vanishes. Only exogenous motion $d_{t}=\|\bar{\Delta}\theta_{t}\|_{g_{\theta_{t}}}$ remains, and its cumulative effect scales as $\sum_{t}d_{t}/T$, as shown in Theorem 13.
Lemma 19 (Performative equilibrium)
If both $d_{t}\equiv 0$ and $\kappa_{t}^{(\mathcal{M})}\equiv 0$, then the process is stationary and $\theta_{t}=\theta_{0}$ for all $t$. The generalization bound collapses to
$\mathbb{E}\,|\hat{R}_{T}-R_{T}|\leq\frac{C_{0}}{\sqrt{T}},$
recovering the standard generalization rate under stationary sampling (Vapnik1998; Perdomo2020).
Proof This is immediate from Theorem 13 by substituting $d_{t}=\kappa_{t}^{(\mathcal{M})}=0$.
Lemma 20 (Adaptive–data regime)
If $d_{t}\equiv 0$ but $\kappa_{t}^{(\mathcal{M})}>0$, then all drift is policy–induced. Theorem 13 becomes
$\mathbb{E}\,|\hat{R}_{T}-R_{T}|\leq\frac{C_{0}}{\sqrt{T}}+\frac{C_{2}}{T}\sum_{t=1}^{T-1}\kappa_{t}^{(\mathcal{M})},$
which parallels the adaptivity penalties in adaptive–data–analysis bounds (Dwork2015).
Proof With $d_{t}=0$, only endogenous drift contributes to nonstationarity. Each step’s deviation $\mathbb{E}_{p_{\theta_{t}}}\ell(f_{t})-\mathbb{E}_{p_{\theta_{t-1}}}\ell(f_{t-1})$ is controlled by Lemma 11 through the Fisher displacement $d_{F}(\theta_{t},\theta_{t-1})$, which reduces here to $\alpha\kappa_{t}^{(\mathcal{M})}$. Summing over $t$ yields the stated bound.
Synthesis.
Lemmas 1720 confirm that Theorem 13 specializes to the familiar regimes under their corresponding assumptions, reducing to
$\mathbb{E}\,|\hat{R}_{T}-R_{T}|=O\!\left(T^{-1/2}+\frac{C_{T}}{T}\right),$
where $C_{T}=\sum_{t}(d_{t}+\alpha\kappa_{t}^{(\mathcal{M})})$ encodes total distributional motion. In other words, nothing essential is lost when we summarize the behavior in terms of the reproducibility budget $C_{T}$.
Appendix B Experiments
All experiments were implemented in Python 3.11 using NumPy, SciPy, and PyTorch (for the nonlinear learner). Per–run integer seeds ensure deterministic reproduction of all reported results. Full source code, configuration files, and figure-generation scripts are available in the public repository (Section C).
Across all settings we record the environment trajectory $\{\theta_{t}\}$, the learner trajectory $\{f_{t}\}$, the empirical trajectory loss
$\hat{R}_{T}\;=\;\frac{1}{T}\sum_{t=1}^{T}\ell(f_{t}(x_{t}),y_{t}),$
and an approximation of the corresponding population loss $R_{T}=\tfrac{1}{T}\sum_{t=1}^{T}R(\theta_{t},f_{t})$, computed either in closed form (linear–Gaussian environment) or using clean mini-batches drawn from the latent generative model (nonlinear environment). The reported metric is the trajectory generalization gap $|\hat{R}_{T}-R_{T}|$. Across all experiments the learner’s training loss decreases substantially over time, ensuring the effects studied arise from drift rather than optimization instability.
B.1 Experiment 1: Stationary Rate and Drift Sensitivity
Objective.
To verify that (i) in the absence of drift the trajectory generalization gap scales as $O(T^{-1/2})$, and (ii) under controlled exogenous or policy-sensitive drift, the gap increases proportionally to the accumulated drift magnitudes.
Setup.
We simulate a linear–Gaussian environment $\theta_{t+1}=A\theta_{t}+Bu_{t}+\eta_{t}$, with $x_{t}\sim\mathcal{N}(\theta_{t},\Sigma)$ and constant Fisher metric $g_{\theta}=\Sigma^{-1}$. Because the dynamics are linear, the drift increments decompose exactly into
$d_{t}\penalty 10000\ =\penalty 10000\ \|\Delta\theta_{t}^{\mathrm{exo}}\|_{g_{\theta}},\qquad\kappa_{t}^{(\mathcal{M})}\penalty 10000\ =\penalty 10000\ \|\Delta\theta_{t}^{\mathrm{endo}}\|_{g_{\theta}}.$
Population risk is available in closed form via $\mathrm{tr}(\Sigma)+\|\theta_{t}-\hat{\mu}_{t}\|_{2}^{2}$.
Findings.
The $\gamma=0$ (i.i.d.) regime recovers the predicted $T^{-1/2}$ scaling. When exogenous or policy-sensitive drift is introduced, generalization gaps increase proportionally to the normalized drift magnitudes $T^{-1}\!\sum_{t}d_{t}$ and $T^{-1}\!\sum_{t}\kappa_{t}^{(\mathcal{M})}$.
B.2 Experiment 2: Additivity and Budget Collapse
Objective.
To test the additive structure of drift contributions and the one-dimensional budget collapse predicted by Corollary 16.
Setup.
Using the same linear–Gaussian simulator, we sweep a grid of exogenous drift amplitudes and feedback gains. For each run we record
$\bar{d}\;=\;\frac{1}{T}\sum_{t=1}^{T}d_{t},\qquad\bar{\kappa}\;=\;\frac{1}{T}\sum_{t=1}^{T}\kappa_{t}^{(\mathcal{M})},$
and fit the regression model
$|\hat{R}_{T}-R_{T}|=b_{0}+b_{s}T^{-1/2}+b_{1}\bar{d}+b_{2}\bar{\kappa}.$
The ratio $\alpha_{\star}=b_{2}/b_{1}$ defines an empirical combined budget $C_{T}=\sum_{t}(d_{t}+\alpha_{\star}\kappa_{t}^{(\mathcal{M})})$.
Findings.
A reduced model using only $C_{T}/T$ matches the full three-term model with minimal loss, confirming the predicted budget collapse.
B.3 Experiment 3: Metric Alignment
Objective.
To isolate the geometric role of optimization metric alignment.
Setup.
We compare Euclidean and natural-gradient updates that achieve the same multiplicative decrease in a Euclidean quadratic objective $J(\theta)=\frac{1}{2}\|\theta\|_{2}^{2}$, under a fixed Fisher metric $g_{\theta}=\Sigma^{-1}$. Differences in cumulative Fisher path length
$A\;=\;\sum_{t}\|\Delta\theta_{t}\|_{g_{\theta_{t}}}$
reflect only the geometry of the update directions.
Findings.
Natural-gradient trajectories incur substantially shorter Fisher path lengths than Euclidean ones, consistent with interpreting $C_{T}$ as a measure of geometric motion.
B.4 Experiment 4: Nonlinear Speed–Limit Validation
Objective.
To test whether the speed–limit relation persists in a nonlinear learner that violates the simplifying assumptions of the theory.
Latent environment and learner.
Inputs are drawn as $x_{t}\sim\mathcal{N}(0,I_{d})$ and labels as
$y_{t}=\phi(x_{t})^{\top}\theta_{t}+\sigma\xi_{t},\qquad\xi_{t}\sim\mathcal{N}(0,1),$
with $\phi$ a fixed random feature map. A two-layer MLP is trained using single-sample SGD. Population risk is approximated periodically using clean batches drawn from the latent generative model.
Drift construction.
At each step an exogenous direction (unit length in the Fisher metric $G(\theta_{t})$ of the latent model) and a policy-sensitive direction (derived from network disagreement on a probe batch) are proposed. Their expected Fisher lengths per step are $C_{\mathrm{exo}}/T$ and $\gamma/T$; budget managers truncate increments to enforce these rates. We record surrogate drift magnitudes
$d_{t}=\|\Delta\theta_{t}^{\mathrm{exo}}\|_{G(\theta_{t})},\qquad\kappa_{t}^{(\mathcal{M})}=\|\Delta\theta_{t}^{\mathrm{endo}}\|_{G(\theta_{t})}.$
Calibration and validation.
Calibration experiments fit
$|\hat{R}_{T}-R_{T}|=b_{0}+b_{s}T^{-1/2}+b_{1}\bar{d}+b_{2}\bar{\kappa},$
yielding $\alpha_{\star}=b_{2}/b_{1}$. Held-out runs at larger horizons test the reduced model
$|\hat{R}_{T}-R_{T}|=c_{0}+c_{1}T^{-1/2}+c_{2}(C_{T}/T).$
The reduced model accurately predicts the empirical gaps despite the nonlinear learner dynamics.
Appendix C Reproducibility Artifacts
A complete repository containing source code, configuration files, and figure-generation scripts is available at:
https://github.com/fiazaich/Learning-under-Drift